### v1

# Graphical Analysis of Nonlinear Multivariable Feedback Systems

Abstract
Scaled Relative Graphs (SRGs) provide a novel graphical frequency-domain method for the analysis of nonlinear systems. There have been recent efforts to generalize SRG analysis to Multiple-Input Multiple-Output (MIMO) systems. However, these attempts yielded only results for square systems, and in some cases, only methods applicable for Linear Time-Invariant (LTI) systems. In this paper, we develop a complete SRG framework for the analysis of MIMO systems, which may be nonlinear and non-square. The key element is the embedding of operators to a space of operators acting on a common Hilbert space, while restricting the input space to the original input dimension. We develop interconnection rules that use restricted input spaces and stability theorems to guarantee causality, well-posedness and (incremental) $L_{2}$-gain bounds for the overall interconnection. We show utilization of the proposed theoretical concepts on the analysis of nonlinear systems in a linear fractional representation form, which is a rather general class of systems with a representation form directly utilizable for control. Moreover, we provide formulas for the computation of MIMO SRGs of stable LTI operators and diagonal static nonlinear operators. Finally, we demonstrate the capabilities of our proposed approach on several examples.
{IEEEkeywords}
Nonlinear systems, Multivariable systems, Scaled Relative Graph, Incremental stability, Robust stability, Stability of nonlinear systems
1 Introduction
[…]
Graphical methods, such as the Nyquist [1] and Bode [2] diagrams, are foundational to control engineering. They offer intuitive, visually interpretable tools for assessing stability and performance directly from frequency-response data, and provide necessary and sufficient conditions for stability. Techniques ranging from manual loop shaping, sequential loop closing, to controller autotuning rely on these methods, which are central to modern control design. Moreover, the gain and phase margins from the Nyquist diagram have been core building blocks that led to powerful methods of modern robust control [3]. These tools unify the analysis of well-posedness, stability, and performance for Linear Time-Invariant (LTI) multivariable feedback systems with arbitrary number of inputs and outputs, even under uncertainty, and enable optimal controller synthesis both in the model-based based and data-driven cases. However, the LTI property is a severe limitation for the applicability of these tools and extending them to multivariable nonlinear systems with arbitrary number of inputs and outputs remains an open challenge.
The Scaled Relative Graph (SRG) [4] was recently proposed as a novel graphical framework for analyzing nonlinear Single-Input Single-Output (SISO) systems [5]. The SRG offers a non-approximative method yielding sufficient conditions for stability and upper bounds on the (incremental) $L_{2}$-gain, a key performance metric in practice. It is modular, allowing interconnections to be analyzed by composing the SRGs of subsystems. The SRG recovers classical results such as the small-gain theorem and generalizes the circle criterion [6], due to its close relation to the Nyquist diagram. It also enables frequency-dependent gain bounds, forming the basis of a nonlinear Bode diagram and bandwidth definition [7], and accommodates for the LTI notions of phase lead/lag [8]. Apart from theoretical developments, SRGs have proven effective in applications such as reset control analysis [9] and design [10], and circuit modeling [11, 12].
There also have been multiple efforts to extend the SISO SRG tools to the Multiple-Input Multiple-Output (MIMO) setting, but only to a limited degree. They either consider only square [^1] Systems with the same amount of inputs as outputs. LTI systems [13, 14, 15, 16], or when they do allow nonlinear systems, but they remain limited to square systems [17, 18]. In the presence of nonlinearities, well-posedness of the feedback system is assumed in [18]. This assumption provides serious limitation, since well-posedness is easy to violate in the MIMO setting due to interactions among the input and output channels, and its satisfaction is a vital part of multivariable system analysis methods such as the robust control framework. [3].
In essence, we recognize that there exists no modular framework that can verify and quantify well-posedness, stability and performance of interconnections of MIMO operators, either LTI, nonlinear and/or non-square. Even if the SRG framework promises a new avenue for computational nonlinear system analysis, its applicability for systems with arbitrary numbers of inputs and outputs is still lacking, excluding many practically relevant MIMO applications where input-output interaction is a key feature of the dynamics. Development of a framework to handle such systems is thus essential for broader applicability in real-world control scenarios.
In this paper, we introduce a framework for analyzing nonlinear MIMO feedback systems that provides well-posedness, stability and (incremental) $L_{2}$-gain performance bounds, where SRGs play the role of the graphical computational tool. The resulting framework is non-approximative, in contrast to using robust control tools [3] where nonlinearities are embedded into linear uncertainties.
The main obstruction to generalizing the SRG as introduced in [5] to MIMO systems is the fact that the SRG is defined for operators that act on a Hilbert space. When operators are not square, the Hilbert space of inputs is not the same as the space of outputs, and artificially adding inputs/outputs to make the system square leads to conservatism. The key idea of this paper is to embed all operators in an interconnection to the same space of operators on a Hilbert space and carry out the SRG interconnections in a common Hilbert space. Conservatism due to the additional input/output dimensions is removed by restricting the space of input signals to the original input dimension. We state and prove interconnection rules that use restricted input spaces, both for the incremental and non-incremental case. Using these interconnection rules, we provide stability theorems for nonlinear systems that guarantee well-posedness, causality and incremental $L_{2}$-gain bounds, which is achieved by generalizing the (incremental) homotopy results from [19] to the non-square MIMO setting. Throughout the paper, systems in Linear Fractional Representation (LFR) form are considered, i.e. systems that are composed of an LTI block connected to a (diagonal) block of static nonlinearities via fractional transformation, to show applicability of the proposed theoretical concepts. Such LFR representations of nonlinear systems are used thoroughly in robust control [3], but have also been proven effective for modeling [20, 21] and identifying [22, 23, 24] nonlinear systems due to their ability to express a broad range of dynamical phenomena. For stable LTI systems, which need not be square, we provide a computationally tractable algorithm to bound its SRG, composed of computations using singular value decomposition. We also provide bounds of the SRG of a diagonal static nonlinear block. We demonstrate our methods on three examples; a SISO system with multiple nonlinearities, a MIMO mass-spring-damper system with multiple nonlinear springs, and the example in [18]. On the latter, we compare our methods with the approach proposed in [18].
Our theory comes with a software toolbox in Julia for all the necessary computations for SRG based analysis and performance analysis of non-square MIMO nonlinear systems. Using the toolbox, we have created scripts to generate all figures and results in this paper. The toolbox and scripts are freely available at github.com/Krebbekx/SrgTools.jl.
The paper is structured as follows. In Section 2, we present the required preliminaries that lay out the mathematical setting. In Section 3, we introduce systems in LFR form as the canonical interconnection we study in this paper. The SRG of MIMO operators is defined in Section 4, where the embedding procedure is detailed. Next, we develop interconnection rules for MIMO operators in Section 5 and apply it to the LFR. In Section 6, we state and prove our stability theorems for system analysis, which includes a practical theorem for systems in LFR form. We provide formulas to bound the MIMO SRG of stable LTI operators and common diagonal nonlinear operators in Section 7. We demonstrate our methods on three examples in Section 8 and present our conclusions in Section 9.
2 Preliminaries
2.1 Notation and Conventions
[…]
2.2 Operators, Signals, Systems and Stability
2.2.1 Relations and Operators
A relation $R:X\to Y$ is a possibly multi-valued map, defined by $Rx\subseteq Y$ for all $x\in X=:\mathrm{dom}(R)$, and the range is defined as $\mathrm{ran}(R):=\{y\in Y\mid\exists x\in X:y\in Rx\}\subseteq Y$. The graph of a relation $R$ is the set $\{(x,y)\in X\times Y\mid x\in X,y\in Rx\}$. Given the sets $X,Y,Z$ and relations $R,S:X\to Y$ and $T:Y\to Z$, the inverse $R^{-1}$, sum $R+S$ and product $TR$ are defined as
$R^{-1}$ =$=\{(y,x)\mid(x,y)\in R\},$ (1a)
$R+S$ =$=\{(x,y+z)\mid(x,y)\in R,(x,z)\in S\},$ (1b)
$TR$ =$=\{(x,z)\mid\exists y:(x,y)\in R,\,(y,z)\in T\}.$ (1c)
A single-valued relation is called an operator, where $Rx\in Y$ is understood. We denote the set of operators from $X$ to $Y$ as
$\mathscr{N}(X,Y):=\{R:X\to Y\mid\forall x\in X,\,Rx\in Y\},$ (2)
where $\mathscr{N}(X,X)=:\mathscr{N}(X)$. For an operator $R\in\mathscr{N}(X,Y)$, we denote the domain as $\mathrm{dom}(R)=X$ and the range $\mathrm{ran}(R)=\{y\in Y\mid\exists x\in X:y=Rx\}\subseteq Y$. The identity operator on a space $X$ is defined as $I_{X}x=x,\,\forall x\in X$. If $R\in\mathscr{N}(X,Y)$ is injective, then $R^{-1}\in\mathscr{N}(\mathrm{ran}(R),X)$ such that $R^{-1}R=I_{X}$ and $RR^{-1}=I_{\mathrm{ran}(R)}$. If $R$ is not injective, then $R^{-1}$ is multivalued.
For an operator $R:X\to Y$, where $(X,\left\lVert\cdot\right\rVert_{X})$ and $(Y,\left\lVert\cdot\right\rVert_{Y})$ are normed spaces, we define the induced incremental norm as (similar to the notation in [25])
$\Gamma(R):=\sup_{x_{1},x_{2}\in X}\frac{\left\lVert Rx_{1}-Rx_{2}\right\rVert_{Y}}{\left\lVert x_{1}-x_{2}\right\rVert_{X}}\in[0,\infty].$ (3)
Similarly, we define the induced non-incremental norm as
$\gamma(R):=\sup_{x\in X}\frac{\left\lVert Rx\right\rVert_{Y}}{\left\lVert x\right\rVert_{X}}\in[0,\infty].$ (4)
Since the Banach fixed point theorem plays an important role in the development of some results, we state it in detail.
Theorem 1 (Thm. 5.1.1 [26]).
Let $X$ be a non-empty complete metric space and $T:X\to X$ such that for some $10<L<$ 1
$\left\lVert Tx-Ty\right\rVert\leq L\left\lVert x-y\right\rVert,\quad\forall x,y\in X.$
Then, there exists a unique $x^{*}\in X$ such that $Tx^{*}=x^{*}$.
Note that if $X$ is a Banach space and $\mathscr{U}\subseteq X$ is closed, then $\mathscr{U}$ is a complete metric space [27].
2.2.2 Signal Spaces
Let $\mathscr{L}$ denote a Hilbert space, equipped with an inner product $\left\langle\cdot,\cdot\right\rangle_{\mathscr{L}}:\mathscr{L}\times\mathscr{L}\to\mathbb{C}$ and norm $\left\lVert x\right\rVert_{\mathscr{L}}:=\sqrt{\left\langle x,x\right\rangle_{\mathscr{L}}}$. For $d\in\mathbb{N}^{+}$, $\mathbb{T}\in\{[0,T],\mathbb{R}_{\geq 0}\mid T>0\}$ and $\mathbb{F}\in\{\mathbb{R},\mathbb{C}\}$, the Hilbert spaces of interest are
$L_{2}^{d}(\mathbb{T},\mathbb{F}):=\{f:\mathbb{T}\to\mathbb{F}^{d}\mid\left\lVert f\right\rVert<\infty\},$ (5)
[…]
An element $f\in L_{2}^{d}(\mathbb{T},\mathbb{F})$ is denoted as $f=(f_{1},\dots,f_{d})^{\top}$, where $f_{i}\in L_{2}(\mathbb{T},\mathbb{F})$ for all $i=1,\dots,d$. The zero element $0\in L_{2}^{d}(\mathbb{T},\mathbb{F})$ refers to the map $\mathbb{T}\ni t\mapsto 0\in\mathbb{F}^{d}$. For $n\leq d$, we define the linear subspaces (which are Banach spaces)
$\mathscr{U}^{d}_{n}:=\{f\in L_{2}^{d}(\mathbb{T},\mathbb{F})\mid f_{i}=0\text{for }i>n\},$ (6)
where the superscript $d$ is dropped if it is clear from the context.
For any $T\in\mathbb{R}_{\geq 0}$, define the truncation operator $P_{T}:L_{2}^{d}(\mathbb{F})\to L_{2}^{d}(\mathbb{F})$ as
$(P_{T}u)(t):=\begin{cases}u(t)&t\leq T,\\ 0&t>T.\end{cases}$
The extension of $L_{2}^{d}(\mathbb{F})$, see Ref. [28], is defined as
$L_{2\mathrm{e}}^{d}(\mathbb{F}):=\{u:\mathbb{R}_{\geq 0}\to\mathbb{F}^{d}\mid\left\lVert P_{T}u\right\rVert_{2}<\infty\text{ for all }T\in\mathbb{R}_{\geq 0}\},$
where we abbreviate $L_{2\mathrm{e}}^{d}(\mathbb{R}_{\geq 0},\mathbb{F})=:L_{2\mathrm{e}}^{d}(\mathbb{F})$ and $L_{2\mathrm{e}}^{d}(\mathbb{R})=:L_{2\mathrm{e}}^{d}$, and the superscript $d$ is dropped if $d=1$. The extended space is the natural setting for modeling systems, as it includes periodic signals, which are otherwise excluded from $L_{2}$. However, extended spaces are not even normed spaces [29, Ch. 2.3]. Therefore, the Hilbert space $L_{2}^{d}(\mathbb{T},\mathbb{F})$ is the adequate signal space for rigorous functional analytic system analysis.
2.2.3 Systems and Stability
An operator $R$ is said to be causal on $L_{2}^{p}$ ($L_{2\mathrm{e}}^{p}$) if it satisfies $R:L_{2}^{p}\to L_{2}^{q}$ ($R:L_{2\mathrm{e}}^{p}\to L_{2\mathrm{e}}^{q}$) and $P_{T}(Ru)=P_{T}(R(P_{T}u))$ for all $u\in L_{2}^{p}$ ($u\in L_{2\mathrm{e}}^{p}$) and $T\in\mathbb{R}_{\geq 0}$, i.e., the output at time $t$ is independent of the signal at times greater than $t$. Unless otherwise specified, we will always assume causality on $L_{2}^{p}$. Causal systems $R:L_{2}^{p}\to L_{2}^{q}$ are extended to $L_{2\mathrm{e}}^{p}$ by defining $R:L_{2\mathrm{e}}^{p}\to L_{2\mathrm{e}}^{q}$ as $P_{T}Ru:=P_{T}RP_{T}u$, which is well-defined since $P_{T}u\in L_{2}^{p}$ for all $u\in L_{2\mathrm{e}}^{p}$. If $R:L_{2}^{p}\to L_{2}^{q}$ and $R:L_{2\mathrm{e}}^{p}\to L_{2\mathrm{e}}^{q}$, then $R$ is causal on $L_{2}^{p}$ if and only if $R$ is causal on $L_{2\mathrm{e}}^{p}$ [29, Ch. 2.4]. If $R:\mathrm{dom}(R)\subsetneq L_{2}^{p}\to L_{2}^{q}$, then $R$ can only be extended to $L_{2\mathrm{e}}^{p}$ if $\left\lVert P_{T}Ru\right\rVert_{2}<\infty$ for all $u\in L_{2}^{p}$ and $T\in\mathbb{R}_{\geq 0}$, i.e. no finite escape time. Conversely, if $R:L_{2\mathrm{e}}^{p}\to L_{2\mathrm{e}}^{q}$, then it can be that $Ru\notin L_{2}^{q}$ for all $u\in L_{2}^{p}$ (e.g., consider $u(t)\mapsto\sin(t)$). We model physical systems as operators that take inputs in an extended signal space, i.e. $R:L_{2\mathrm{e}}^{p}\to L_{2\mathrm{e}}^{q}$, and we always assume $R(0)=0$, unless otherwise specified.
A system $R$ is said to be $L_{2}$-stable if $R:L_{2}^{p}\to L_{2}^{q}$. For an $L_{2}$-stable system, we define the (non-)incremental $L_{2}$-gain as $\Gamma(R)$ ($\gamma(R)$) from Eq. (3) (Eq. (4)). When $R:L_{2}^{p}\to L_{2}^{q}$ and $\Gamma(R)<\infty$ ($\gamma(R)<\infty$), we call the system (non-)incrementally stable. The general approach in this work is to show that $\Gamma(R)<\infty$ ($\gamma(R)<\infty$) on $\mathrm{dom}(R)\subseteq L_{2}^{p}$ and separately show that $\mathrm{dom}(R)=L_{2}^{p}$. The final step is to extend the domain to $L_{2\mathrm{e}}^{p}$ by proving, or assuming, causality.
Note that the (non-)incremental gain $\Gamma(R)$ ($\gamma(R)$) is computed using signals in $L_{2}^{p}$ only. If $R:L_{2}^{p}\to L_{2}^{q}$ happens to be causal, this gain carries over to $L_{2\mathrm{e}}^{p}$ in the sense that [6]
$\Gamma(R)=\sup_{u_{1},u_{2}\in L_{2\mathrm{e}}^{p}}\sup_{T\in\mathbb{R}_{\geq 0}}\frac{\left\lVert P_{T}Ru_{1}-P_{T}Ru_{2}\right\rVert}{\left\lVert P_{T}u_{1}-P_{T}u_{2}\right\rVert},$
and similarly for $\gamma(R)$.
2.3 The Scaled Relative Graph
Let $\mathscr{L}$ be a Hilbert space, and $R:\mathrm{dom}(R)\subseteq\mathscr{L}\to\mathscr{L}$ a relation. Define the angle between $u,y\in\mathscr{L}$ as
$\angle(u,y):=\cos^{-1}\frac{\mathrm{Re}\left\langle u,y\right\rangle}{\left\lVert u\right\rVert\left\lVert y\right\rVert}\in[0,\pi].$ (7)
Given distinct $u_{1},u_{2}\in\mathscr{U}\subseteq\mathrm{dom}(R)$, we define the set
$z_{R}(u_{1},u_{2}):=\\ \left\{\frac{\left\lVert y_{1}-y_{2}\right\rVert}{\left\lVert u_{1}-u_{2}\right\rVert}e^{\pm j\angle(u_{1}-u_{2},y_{1}-y_{2})}\mid y_{1}\in Ru_{1},y_{2}\in Ru_{2}\right\}\\ \cup\{\infty\mid\text{if }R\text{ is multi-valued }\}.$
The SRG of $R$ over the set $\mathscr{U}$ is defined as
$\operatorname{SRG}_{\mathscr{U}}(R):=\bigcup_{u_{1},u_{2}\in\mathscr{U},\;u_{1}\neq u_{2}}z_{R}(u_{1},u_{2})\subseteq\mathbb{C}_{\infty},$
and we denote $\operatorname{SRG}(R):=\operatorname{SRG}_{\mathrm{dom}(R)}(R)$. Note that $0\in\operatorname{SRG}_{\mathscr{U}}(R)$, if and only if there exist $u_{1},u_{2}\in\mathscr{U}$, $u_{1}\neq u_{2}$, such that $Ru_{1}=Ru_{2}$.
One can also define the Scaled Graph (SG) around some particular input. The SG of an operator $R$ with one input $u^{\star}\in\mathrm{dom}(R)$ fixed and the other in set $\mathscr{U}$ is defined as
$\operatorname{SG}_{\mathscr{U},u^{\star}}(R):=\{z_{R}(u,u^{\star})\mid u\in\mathscr{U}\setminus u^{\star}\}.$ (8)
Again, we use the shorthand $\operatorname{SG}_{\mathrm{dom}(R),u^{\star}}(R)=\operatorname{SG}_{u^{\star}}(R)$. The SG around $u^{\star}=0$ is particularly interesting, because the radius of $\operatorname{SG}_{0}(R)$ gives a non-incremental $L_{2}$-gain bound for the operator $R$. By definition of the SRG (SG), the (non-)incremental gain of an operator $R:\mathscr{L}\to\mathscr{L}$, defined in Eq. (3) (Eq. (4)), is equal to the radius of the SRG (SG at zero), i.e. $\Gamma(R)=r_{\mathrm{min}}(\operatorname{SRG}(R))$ ($\gamma(R)=r_{\mathrm{min}}(\operatorname{SG}_{0}(R))$).
For parallel and series interconnections, we require the definitions of chords and arcs [4]. Denote the line segment between $z_{1},z_{2}\in\mathbb{C}$ as $[z_{1},z_{2}]:=\{\alpha z_{1}+(1-\alpha)z_{2}\mid\alpha\in[0,1]\}$. Let the right-hand arc, denoted by $\operatorname{Arc}^{+}(z,\bar{z})$, be the circle segment of the circle that is centered at the origin and intersects $z,\bar{z}$, with real part greater than $\mathrm{Re}(z)$. The left-hand arc, denoted by $\operatorname{Arc}^{-}(z,\bar{z})$, is similarly defined, but with real part smaller than $\mathrm{Re}(z)$. More precisely
$\operatorname{Arc}^{+}(z,\bar{z})=$$\{re^{j(1-2\alpha)\phi}$
$\,\mid z=re^{j\phi},\phi\in(-\pi,\pi],\alpha\in[0,1]\},$
$\operatorname{Arc}^{-}(z,\bar{z})=$ $-\operatorname{Arc}^{+}(-z,-\bar{z}).$
Definition 1.
A set $\mathscr{C}\subseteq\mathbb{C}$ is said to satisfy the chord property if for all $z\in\mathbb{C}$, it holds that $[z,\bar{z}]\subseteq\mathscr{C}$.
Definition 2.
A set $\mathscr{C}\subseteq\mathbb{C}$ is said to satisfy the left-arc (right-arc) property if for all $z\in\mathbb{C}$, it holds that $\operatorname{Arc}^{-}(z,\bar{z})\subseteq\mathscr{C}$ ($\operatorname{Arc}^{+}(z,\bar{z})\subseteq\mathscr{C}$). If $\mathscr{C}$ satisfies the left-arc and/or right-arc property, it is said to satisfy an arc property.
3 Nonlinear Systems in LFR Form
[…]
(a) A general feedback interconnection from [28].
$G$$\Phi$$y$$u$$z$$w$
(b) A system in LFR form.
Figure 1: Two canonical feedback interconnections.
The analysis of nonlinear feedback systems from the input/output perspective often considers the system depicted in Fig. 1(a). This interconnection, where $H_{1}$ and $H_{2}$ are nonlinear operators, is the canonical system in [29, 28] and is used to develop all classical results: the small gain and passivity theorems, and in the case that $H_{1}$ is LTI, the circle and Popov criteria. Even though the system in Fig. 1(a) covers many feedback systems, there are many situations that require a more sophisticated modeling approach.
Instead, we will consider systems in LFR form, as shown in Fig. 1(b), where $G$ contains all LTI dynamics and $\Phi$ all static or dynamic nonlinearities. The LFR is used thoroughly in the robust control framework [3] for uncertain LTI systems, for nonlinear system identification [22, 23, 24], to machine learning [20], and for analysis of recurrent neural networks [21]. The reason for this is that a very broad class of systems can be written in LFR form. For example, the system in Fig. 1(a) can always be represented in LFR form, and it is even true that for most dynamic nonlinearities, by appropriate loop transformation, all dynamics can be gathered in the LTI block $G$, while $\Phi$ becomes static diagonal. Therefore, we will focus on systems in LFR form as the leading example throughout this paper. We emphasize, however, that the applicability of the framework developed in this paper reaches beyond the feedback systems in Fig. 1.
3.1 Definition of Systems in LFR Form
A system $R:L_{2}^{p}\to L_{2}^{q}$ is said to be in LFR form if it is decomposed as the interconnection in Fig. 1(b), by considering the partition of the LTI operator $G\in RH_{\infty}^{(q+n_{\mathrm{z}})\times(p+n_{\mathrm{w}})}$
$G=\begin{pmatrix}G_{\mathrm{zw}}&G_{\mathrm{zu}}\\ G_{\mathrm{yw}}&G_{\mathrm{yu}}\end{pmatrix},$ (9)
where $G_{\mathrm{zw}}\in RH_{\infty}^{n_{\mathrm{z}}\times n_{\mathrm{w}}}$, $G_{\mathrm{zu}}\in RH_{\infty}^{n_{\mathrm{z}}\times p}$, $G_{\mathrm{yw}}\in RH_{\infty}^{q\times n_{\mathrm{w}}}$ and $G_{\mathrm{yu}}\in RH_{\infty}^{q\times p}$, and a nonlinear operator
$\Phi:L_{2}^{n_{\mathrm{z}}}\to L_{2}^{n_{\mathrm{w}}}.$ (10)
The LTI block $G$ represents the LTI dynamics, and all static or dynamic nonlinearities are collected in $\Phi$. Note that $\Phi$ may also be used to represent uncertain nonlinear effects. In the robust control framework [3], one collects all linear uncertainties in $\Phi$, commonly denoted as $\Delta$ instead.
By connecting $\Phi$ to the $w$ and $z$ channels as shown in Fig. 1(b), one obtains the closed-loop operator $y=Ru$ as
$R$ $=G_{\mathrm{yw}}\Phi(I-G_{\mathrm{zw}}\Phi)^{-1}G_{\mathrm{zu}}+G_{\mathrm{yu}}$
$=G_{\mathrm{yw}}(\Phi^{-1}-G_{\mathrm{zw}})^{-1}G_{\mathrm{zu}}+G_{\mathrm{yu}}.$ (11)
The main objective in this paper is to analyze the stability, well-posedness and $L_{2}$-gain of the closed-loop system in (11). In the SISO case, the strategy to analyze Eq. (11) using SRG methods would be to replace each individual operator with its SRG, and resolve the sum, product and inverse relations using the so-called SRG calculus [6]. However, the dimensions $p,q,n_{\mathrm{w}},n_{\mathrm{z}}$ may take any value in $\mathbb{N}_{+}$, so the operators in Eq. (11) can be non-square. Therefore, the LFR form exemplifies the need for an SRG framework that can deal with non-square MIMO operators.
4 The SRG for MIMO Operators
At the core of the SRG definition lies the existence of a Hilbert space that contains both the domain and range. For an operator $R:L_{2}^{p}\to L_{2}^{q}$, this definition is problematic when $p\neq q$, i.e. the operator is not square. As a first step towards a general framework for SRG analysis of multivariable feedback systems, we will develop the required mathematical tools to embed operators $R\in\mathscr{N}(L_{2}^{p},L_{2}^{q})$ into $\mathscr{N}(L_{2}^{n})$ for some $n\geq p,q$. Using these embeddings, we will define the SRG for a general MIMO operator, called the MIMO SRG.
4.1 Embedding of MIMO Operators
Throughout this section, we consider $p,q\in\mathbb{N}_{+}$ and $R:L_{2}^{p}\to L_{2}^{q}$, and define $n:=\max\{p,q\}$. One can be in one of three cases: $p>q$ (flat), $p<q$ (tall) and $p=q$ (square). For each case, we will consider the appropriate embedding.
4.1.1 Flat operators
When $p>q$, we append outputs that map to zero by defining the linear embedding operator
$\iota_{p\leftarrow q}$ $:L_{2}^{q}\to L_{2}^{p},$ (12)
$(f_{1},\dots,f_{q})^{\top}$ $\mapsto(f_{1},\dots,f_{q},\underbrace{0\dots,0}_{p-q})^{\top}.$
This way, $\iota_{p\leftarrow q}$ embeds $L_{2}^{p}$ into $L_{2}^{q}$ and $R\mapsto\iota_{p\leftarrow q}R$ embeds $\mathscr{N}(L_{2}^{p},L_{2}^{q})$ into $\mathscr{N}(L_{2}^{p})$.
Lemma 1.
The embedding operator $\iota_{p\leftarrow q}:L_{2}^{q}\to\mathscr{U}_{q}^{p}$ is an isometric isomorphism.
The proof can be found in the Appendix.
For a flat operator $R$ and $\mathscr{U}\subseteq L_{2}^{p}$, we define the SRG as
$\operatorname{SRG}_{\mathscr{U}}(R):=\operatorname{SRG}_{\mathscr{U}}(\iota_{p\leftarrow q}R).$ (13)
It is important to note that $\frac{\left\lVert Ru\right\rVert}{\left\lVert u\right\rVert}=\frac{\left\lVert\iota_{p\leftarrow q}Ru\right\rVert}{\left\lVert u\right\rVert}$ for all $u\in L_{2}^{p}$, which guarantees the important relations $\Gamma(R)=r_{\mathrm{min}}(\operatorname{SRG}(R))$ and $\gamma(R)=r_{\mathrm{min}}(\operatorname{SG}_{0}(R))$.
4.1.2 Tall operators
When $p<q$, we add extra inputs that are ignored in the output. For this, we define the linear projection operator
$\pi_{p\leftarrow q}$ $:L_{2}^{q}\to L_{2}^{p},$ (14)
$(f_{1},\dots,f_{p},r_{1},\dots,r_{q-p})^{\top}$ $\mapsto(f_{1},\dots,f_{p})^{\top}.$
This way, $\pi_{p\leftarrow q}$ projects $L_{2}^{q}$ onto $L_{2}^{p}$ and $R\mapsto R\pi_{p\leftarrow q}$ embeds $\mathscr{N}(L_{2}^{p},L_{2}^{q})$ into $\mathscr{N}(L_{2}^{q})$.
Lemma 2.
The projection operator $\pi_{p\leftarrow q}:\mathscr{U}_{p}^{q}\to L_{2}^{p}$ is an isometric isomorphism with inverse $\pi_{p\leftarrow q}^{-1}=\iota_{q\leftarrow p}$
The proof can be found in the Appendix.
The extra inputs $r_{1},\dots,r_{q-p}$ should not belong to the description of $R$. Therefore, we restrict the inputs of $R\pi_{p\leftarrow q}$ to the space $\mathscr{U}_{p}^{q}$ in Eq. (6). For a tall operator $R$ and $\mathscr{U}\subseteq L_{2}^{p}$, we define the SRG as
$\operatorname{SRG}_{\mathscr{U}}(R):=\operatorname{SRG}_{\iota_{q\leftarrow p}(\mathscr{U})}(R\pi_{p\leftarrow q}).$ (15)
Note that $u=\pi_{p\leftarrow q}\iota_{q\leftarrow p}u$, so $\iota_{q\leftarrow p}(\mathscr{U})$ precisely contains all inputs to $R$. Therefore, we can conclude $\frac{\left\lVert Ru\right\rVert}{\left\lVert u\right\rVert}=\frac{\left\lVert Ru\right\rVert}{\left\lVert\iota_{q\leftarrow p}u\right\rVert}$ for all $u\in L_{2}^{p}$, which guarantees the important relations $\Gamma(R)=r_{\mathrm{min}}(\operatorname{SRG}(R))$ and $\gamma(R)=r_{\mathrm{min}}(\operatorname{SG}_{0}(R))$.
4.1.3 Square operators
Now consider $R:L_{2}^{n}\to L_{2}^{n}$ and choose any $\tilde{n}>n$. We embed $R$ into $\mathscr{N}(L_{2}^{\tilde{n}})$ using the map $R\mapsto\iota_{\tilde{n}\leftarrow n}R\pi_{n\leftarrow\tilde{n}}$. Using the appropriate input space, this embedding can be done without introducing conservatism.
Lemma 3.
Let $R:L_{2}^{n}\to L_{2}^{n}$ and choose any $\tilde{n}>n$, then
$\operatorname{SRG}_{\mathscr{U}}(R)=\operatorname{SRG}_{\iota_{\tilde{n}\leftarrow n}(\mathscr{U})}(\iota_{\tilde{n}\leftarrow n}R\pi_{n\leftarrow\tilde{n}}).$ (16)
The proof can be found in the Appendix.
4.2 Definition of the MIMO SRG
We are now ready to define the SRG for a general operator $R:L_{2}^{p}\to L_{2}^{q}$.
Definition 3.
The SRG of an operator $R:L_{2}^{p}\to L_{2}^{q}$ over $\mathscr{U}\subseteq L_{2}^{p}$ is defined as
$\operatorname{SRG}_{\mathscr{U}}(R):=\operatorname{SRG}_{\iota_{n\leftarrow p}(\mathscr{U})}(\iota_{n\leftarrow q}R\pi_{p\leftarrow n}),$ (17)
where $n\geq\max\{p,q\}$.
Note that by Lemma 3, the choice of $n\geq\max\{p,q\}$ in Definition 3 does not affect the MIMO SRG. A key feature of Definition 3 is that the embedding of the domain $\iota_{n\leftarrow p}(\mathscr{U})$ does not artificially introduce extra inputs to the embedded operator $\iota_{n\leftarrow q}R\pi_{p\leftarrow n}$, which are otherwise not present in the description of $R$. This allows operators to be embedded in the same space of square operators without introducing any conservatism in the embedding step.
5 Interconnecting MIMO Operators
Using the MIMO SRG from Definition 3, we have the tools to define the SRG of each operator of the LFR in Eq. (11). The next step is to derive formulas to study interconnections of these operators, i.e. operator inverses, sums and compositions. Formulas that bound the SRG of operator inverses, sums and compositions are derived in [4]. However, little attention is given to the domain and range of the operators under consideration, which can lead to errors as pointed out in [6]. Additionally, the domain of the embedded operator plays a crucial role in the description of the original operator as explained in Section 4.2.
Therefore, we will develop the necessary “calculus” for the SRG and SG, as defined in Section 2.3, that carefully handles the domain and range of operators. We also discuss the notion of adding chords and arcs to an SRG bound, which are necessary for parallel and series connections, respectively. Then, we show how these general rules for SRGs are used to study interconnections of operators using the MIMO SRG from Definition 3, focusing on systems in LFR form. The main result of this section is that once the input/output dimensions match when interconnecting operators, the user does not have to keep track of the technicalities of the embedding in Section 4, and one can analyze interconnections of MIMO operators using the MIMO SRG from Definition 3 and the familiar SRG calculus from [4].
5.1 Interconnection Rules for the SRG
Before we state and prove our interconnection theorems, we discuss how the domain and range of an operator influence the interconnection rules. The five operations we consider are: 1) multiplication with a nonzero real constant, 2) addition with identity, 3) inversion, 4) parallel interconnection and 5) series interconnection. The mathematical definition and their effect on the domain and range are listed in Table 1.
As Table 1 demonstrates, the domain and range of an operator play a nontrivial role during interconnections. For example when connecting two operators $R:\mathrm{dom}(R)\to X,S:\mathrm{dom}(S)\to X$ in parallel, i.e. $R+S$, one may ask the following question: What domain of inputs is contained in $\operatorname{SRG}(R+S)$, and does $\operatorname{SRG}(R)+\operatorname{SRG}(S)$ (analogous to [4]) include all relevant inputs? To address this question, we formulate an SRG interconnection theorem with explicit dependence on the domain and range. The following theorem is a generalization of [4, Section 4].
Theorem 2.
Let $0\neq\alpha\in\mathbb{R}$, let $R,S:X\to Y$, $T:Y\to Z$ be relations on Hilbert spaces $X,Y,Z$ and linear subspaces $\mathscr{U}\subseteq X,\mathscr{Y}\subseteq Y$ such that $R(\mathscr{U})\subseteq\mathscr{Y}$. Then,
  1. a.
    $\operatorname{SRG}_{\mathscr{U}}(\alpha R)=\operatorname{SRG}_{\mathscr{U}}(R\alpha)=\alpha\operatorname{SRG}_{\mathscr{U}}(R)$,
  2. b.
    $\operatorname{SRG}_{\mathscr{U}}(I_{\mathscr{U}}+R)=1+\operatorname{SRG}_{\mathscr{U}}(R)$, where $I_{\mathscr{U}}$ obeys $I_{\mathscr{U}}u=u$ for all $u\in\mathscr{U}$,
  3. c.
    $\operatorname{SRG}_{R(\mathscr{U})}(R^{-1})=(\operatorname{SRG}_{\mathscr{U}}(R))^{-1}$ (where $0,\infty\in\operatorname{SRG}_{\mathscr{U}}(R)$ are allowed).
  4. d.
    If at least one of $\operatorname{SRG}_{\mathscr{U}}(R),\operatorname{SRG}_{\mathscr{U}}(S)$ satisfies the chord property, then $\operatorname{SRG}_{\mathscr{U}}(R+S)\subseteq\operatorname{SRG}_{\mathscr{U}}(R)+\operatorname{SRG}_{\mathscr{U}}(S)$.
  5. e.
    If at least one of $\operatorname{SRG}_{\mathscr{U}}(R),\operatorname{SRG}_{\mathscr{Y}}(T)$ satisfies an arc property, then $\operatorname{SRG}_{\mathscr{U}}(TR)\subseteq\operatorname{SRG}_{\mathscr{Y}}(T)\operatorname{SRG}_{\mathscr{U}}(R)$.
[…]
The proof can be found in the Appendix.
Table 1: Operations and their effect on the domain and range on relations $R,S$ ($0\neq\alpha\in\mathbb{R}$).
Operation Domain Range
$\alpha R$ / $R\alpha$ invariant / $(1/a)\mathrm{dom}(R)$ $\alpha\,\mathrm{ran}(R)$ / invariant
$I+R$ $\mathrm{dom}(I+R)=\mathrm{dom}(R)$ $\mathrm{ran}(I+R)\subseteq$
$\mathrm{ran}(R)+\mathrm{ran}(I)$
$R^{-1}$ $\mathrm{ran}(R)$ $\mathrm{dom}(R)$
$R+S$ $\mathrm{dom}(R+S)=$ $\mathrm{ran}(R+S)\subseteq$
$\mathrm{dom}(R)\cap\mathrm{dom}(S)$ $\mathrm{ran}(R)+\mathrm{ran}(S)$
$RS$ $\mathrm{dom}(RS)\subseteq\mathrm{dom}(S)$ $\mathrm{ran}(RS)\subseteq\mathrm{ran}(R)$
Remark 1.
If $\mathscr{U}$ is just a set and not a linear subspace, the only thing that changes is Theorem 2.a., which becomes $\operatorname{SRG}_{\mathscr{U}}(\alpha R)=\operatorname{SRG}_{(1/a)\mathscr{U}}(R\alpha)=\alpha\operatorname{SRG}_{\mathscr{U}}(R)$, see Table 1.
5.2 Interconnection Rules for the SG
Upon minor modifications, Theorem 2 can be restated for the SG, which is useful for studying non-incremental stability as opposed to incremental stability. The important novel aspect to keep track of in this case is where $u^{\star}$ in $\operatorname{SG}_{\mathscr{U},u^{\star}}(R)$ is mapped to under $R$.
Theorem 3.
Let $0\neq\alpha\in\mathbb{R}$, let $R:X\to Y$ be an operator and $S:X\to Y$, $T:Y\to Z$ be relations on Hilbert spaces $X,Y,Z$ and linear subspaces $\mathscr{U}\subseteq X,\mathscr{Y}\subseteq Y$ such that $R(\mathscr{U})\subseteq\mathscr{Y}$ and $u^{\star}\in X,y^{\star}=Ru^{\star}$. Then,
  1. a.
    $\operatorname{SG}_{\mathscr{U},u^{\star}}(\alpha R)=\alpha\operatorname{SG}_{\mathscr{U},u^{\star}}(R)$ and if $u^{\star}=0$ then also $\operatorname{SG}_{\mathscr{U},u^{\star}}(R\alpha)=\alpha\operatorname{SG}_{\mathscr{U},u^{\star}}(R)$,
  2. b.
    $\operatorname{SG}_{\mathscr{U},u^{\star}}(I_{\mathscr{U}}+R)=1+\operatorname{SRG}_{\mathscr{U}}(R)$, where $I_{\mathscr{U}}$ obeys $I_{\mathscr{U}}u=u$ for all $u\in\mathscr{U}\cup\{u^{\star}\}$,
  3. c.
    $\operatorname{SG}_{R(\mathscr{U}),y^{\star}}(R^{-1})=(\operatorname{SG}_{\mathscr{U},u^{\star}}(R))^{-1}$ (where $0,\infty\in\operatorname{SG}_{\mathscr{U},u^{\star}}(R)$ are allowed).
  4. d.
    If at least one of $\operatorname{SG}_{\mathscr{U},u^{\star}}(R),\operatorname{SG}_{\mathscr{U},u^{\star}}(S)$ satisfies the chord property, then $\operatorname{SG}_{\mathscr{U},u^{\star}}(R+S)\subseteq\operatorname{SG}_{\mathscr{U},u^{\star}}(R)+\operatorname{SG}_{\mathscr{U},u^{\star}}(S)$.
  5. e.
    If at least one of $\operatorname{SG}_{\mathscr{U},u^{\star}}(R),\operatorname{SG}_{\mathscr{Y},y^{\star}}(T)$ satisfies an arc property, then $\operatorname{SG}_{\mathscr{U},u^{\star}}(TR)\subseteq\operatorname{SG}_{\mathscr{Y},y^{\star}}(T)\operatorname{SG}_{\mathscr{U},u^{\star}}(R)$.
[…]
The proof can be found in the Appendix.
Remark 2.
In Theorem 3, we assume that $R$ is single-valued. The theorem can be proven also in the case of relations, but one has to define the SG w.r.t. a set of inputs $u^{\star}\subseteq X$ instead of $u^{\star}\in X$, which is not explored here.
The most frequently used case of Theorem 3 is when $u^{\star}=0$ and $R(u^{\star})=y^{\star}=0$. When $\mathscr{U}=L_{2}$, this situation corresponds to computing the non-incremental gain $\gamma(R)$.
5.3 Adding Chords and Arcs in an Improved Way
When using Theorem 2 (or Theorem 3) to analyze parallel/series interconnections of operators, one must make sure that at least one of the SRG (SG) bounds involved satisfies the chord/arc property, respectively. If this is not the case, one must add chords or arcs to the relevant SRG (SG) bound. In this section, we discuss how to efficently add chords and arcs to complex sets that bound the SRG (or SG) of an operator.
Definition 4.
For $\mathscr{C}\subseteq\mathbb{C}$, define the chord, left-arc (--) and right-arc ($+$) completions, respectively, as
$\mathscr{C}^{\mathrm{c}}:=\bigcup_{z\in\mathscr{C}}[z,\bar{z}],\quad\mathscr{C}^{\mp}:=\bigcup_{z\in\mathscr{C}}\operatorname{Arc}^{\mp}(z,\bar{z}).$
Note that $\mathscr{C}\subseteq\mathscr{C}^{\mathfrak{s}}$, where $\mathfrak{s}\in\{\mathrm{c},-,+\}$.
When taking a sum or product of two relations, which both do not satisfy the chord or arc property, one can take the intersection of all possible completions to obtain an improved (i.e. smaller) bound for the SRG of the sum or product.
Definition 5.
For $\mathscr{C}_{1},\mathscr{C}_{2}\subseteq\mathbb{C}$, the improved chord completion of the sum is defined as
$\overline{\mathscr{C}_{1}+\mathscr{C}_{2}}:=(\mathscr{C}_{1}^{\mathrm{c}}+\mathscr{C}_{2})\cap(\mathscr{C}_{1}+\mathscr{C}_{2}^{\mathrm{c}}).$ (18)
For products, the improved arc completion is defined as
$\overline{\mathscr{C}_{1}\mathscr{C}_{2}}:=(\mathscr{C}_{1}^{+}\mathscr{C}_{2})\cap(\mathscr{C}_{1}\mathscr{C}_{2}^{+})\cap(\mathscr{C}_{1}^{-}\mathscr{C}_{2})\cap(\mathscr{C}_{1}\mathscr{C}_{2}^{-}).$ (19)
The following lemma is useful for bounding the SRG of a sum or product of operators, while adding the least amount of chords/arcs as possible.
Lemma 4.
Let $R,S:X\to Y$, $T:Y\to Z$ be relations on Hilbert spaces $X,Y,Z$ and linear subspaces $\mathscr{U}\subseteq X,\mathscr{Y}\subseteq Y$ such that $R(\mathscr{U})\subseteq\mathscr{Y}$ and $u^{\star}\in X,y^{\star}=Ru^{\star}$, then
$\operatorname{SRG}_{\mathscr{U}}(R+S)\subseteq\overline{\operatorname{SRG}_{\mathscr{U}}(R)+\operatorname{SRG}_{\mathscr{U}}(S)},$
$\operatorname{SRG}_{\mathscr{U}}(TR)\subseteq\overline{\operatorname{SRG}_{\mathscr{Y}}(T)\operatorname{SRG}_{\mathscr{U}}(R)},$
which holds also for $\operatorname{SRG}_{\mathscr{U}}\to\operatorname{SG}_{\mathscr{U},u^{\star}}$, $\operatorname{SRG}_{\mathscr{Y}}\to\operatorname{SG}_{\mathscr{Y},y^{\star}}$.
The proof can be found in the Appendix.
We note that algorithms for performing sums and products with improved chord/arc completions are available at github.com/Krebbekx/SrgTools.jl.
5.4 Interconnecting MIMO Operators
Until now, the results in this section are general, and pertain to the SRG and SG as defined in Section 2.3. We will now discuss how these results can be applied to interconnections of MIMO systems, which may be non-square, using the mathematical tools from Section 4. The SG case, using Theorem 3, is analogous.
Consider the systems $R:L_{2}^{p_{R}}\to L_{2}^{q_{R}},S:L_{2}^{p_{S}}\to L_{2}^{q_{S}}$ and $T:L_{2}^{p_{T}}\to L_{2}^{q_{T}}$, for which we will demonstrate how interconnections are studied. The first step is to compute $n=\max_{i\in\{R,S,T\}}\{p_{i},q_{i}\}$. Then, one computes the MIMO SRG for each operator, as defined in Definition 3. For each operation in Theorem 2, we discuss below how the input and output dimensions influence the SRG analysis of the interconnection.
  1. a.
    Pre/post multiplication with a real gain: For $R$, one has $\mathscr{U}=L_{2}^{p_{R}}$, which is a linear subspace of $\mathscr{U}_{p_{R}}^{n}=\iota_{n\leftarrow p_{R}}(\mathscr{U})$, hence one can apply 2.a. without any further conditions.
  2. b.
    Addition with identity: This operation is well-defined for systems if $p_{R}=q_{R}$. If $p_{R}>q_{R}$, then $I_{\mathscr{U}}$ will have more output dimensions, effectively giving $R$ an extra $p_{R}-q_{R}$ identically zero outputs. Conversely, if $p_{R}<q_{R}$, then the identity only feeds through the first $p_{R}$ inputs and outputs zero for the remaining channels.
  3. c.
    Inversion: This operation is always well-defined. One must keep in mind that $R(L_{2}^{p_{R}})\subsetneq L_{2}^{q_{R}}$ in general. Therefore, if $\operatorname{SRG}(R)^{-1}$ has finite radius, one can only conclude that $R^{-1}$ has finite incremental gain on $\mathrm{dom}(R^{-1})=\mathrm{ran}(R)$, see Table 1.
  4. d.
    Parallel interconnection: Theorem 2.d. assumes a priori that $p_{R}=p_{S}$. For a parallel interconnection to be well-defined, the output dimensions should match as well, i.e. $q_{R}=q_{S}$, which is not enforced by Theorem 2. If (w.l.o.g.) $q_{R}<q_{S}$, then the system $R$ has effectively gained $q_{S}-q_{R}$ identically zero outputs.
  5. e.
    Series interconnection: Theorem 2.e. assumes a priori that $q_{R}\leq p_{T}$ by assuming $\iota_{n\leftarrow q_{R}}R\pi_{p_{R}\leftarrow n}(\mathscr{U}_{p_{R}}^{n})\subseteq\mathscr{U}_{p_{T}}^{n}$. For a series interconnection to be well-defined, the output dimension of $R$ should match the input dimension of $T$, i.e. $q_{R}=p_{T}$. If $q_{R}<p_{T}$, then $R$ has effectively gained $p_{T}-q_{R}$ identically zero output channels. This makes the SRG calculations conservative, since the SRG of $T$ contains more input dimensions than $R$ provides. Conversely, if $q_{R}>p_{T}$, the assumption $R(\mathscr{U})\subseteq\mathscr{Y}$ is violated.
The above shows that, as long as the input/output dimensions match when interconnecting operators, the user can simply use SRG calculus with the symbol $\operatorname{SRG}(\cdot)$, which represents the MIMO SRG in Definition 3. In other words, once the dimensions are right when interconnecting, the subscripts, which indicate the input spaces in Theorems 2 and 3, can be dropped.
Note that if the input/output dimensions do not match, the SRG operations in Theorem 2 are still mathematically well-defined. However, the interconnection of operators as a system has a different interpretation.
To analyze the SRG of an LFR, one first has to compute $n=\max\{p,q,n_{\mathrm{w}},n_{\mathrm{z}}\}$ and embed all operators in Eq. (11) in $\mathscr{N}(L_{2}^{n})$ using Definition 3. Note that for the LFR in Eq. (11), the input/output dimensions match by definition. Therefore, we can use Theorem 2 and Lemma 4 to conclude
$\operatorname{SRG}(R)\subseteq\overline{\operatorname{SRG}(G_{\mathrm{yw}}(\Phi^{-1}-G_{\mathrm{zw}})^{-1}G_{\mathrm{zu}})+\operatorname{SRG}(G_{\mathrm{yu}})},$ (20a)
$\operatorname{SRG}(G_{\mathrm{yw}}(\Phi^{-1}-G_{\mathrm{zw}})^{-1}G_{\mathrm{zu}})\subseteq\\ \overline{\operatorname{SRG}(G_{\mathrm{yw}}(\Phi^{-1}-G_{\mathrm{zw}})^{-1})\operatorname{SRG}(G_{\mathrm{zu}})},$ (20b)
$\operatorname{SRG}(\operatorname{SRG}(G_{\mathrm{yw}}(\Phi^{-1}-G_{\mathrm{zw}})^{-1}))\subseteq\\ \overline{\operatorname{SRG}(G_{\mathrm{yw}})\operatorname{SRG}((\Phi^{-1}-G_{\mathrm{zw}})^{-1})},$ (20c)
$\operatorname{SRG}((\Phi^{-1}-G_{\mathrm{zw}})^{-1})=\operatorname{SRG}(\Phi^{-1}-G_{\mathrm{zw}})^{-1},$ (20d)
$\operatorname{SRG}(\Phi^{-1}-G_{\mathrm{zw}})\subseteq\overline{\operatorname{SRG}(\Phi)^{-1}-\operatorname{SRG}(G_{\mathrm{zw}})}.$ (20e)
Here we used Theorem 2.d. in Eqs. (20a) and (20e), Theorem 2.e. in Eqs. (20b) and (20c) and Theorem 2.c. in Eq. (20d).
We have obtained a bound in Eq. (20) for the SRG of $R$ in Eq. (11), however, we cannot yet conclude stability from this bound alone. This is the topic of the next section.
6 System Analysis using MIMO SRGs
Now, we have developed the tools to describe MIMO operators using SRGs and study interconnections of operators. In particular, we have obtained a bound for the SRG of the LFR in Eq. (11). However, an SRG bound is not enough to characterize the stability, well-posedness and incremental $L_{2}$-gain performance of a feedback system.
The reason for this, as shown in [6], is that inversion of a stable operator $R:L_{2}^{p}\to L_{2}^{q}$ can yield an unstable operator, even if $\operatorname{SRG}(R)^{-1}$ has finite radius. This can be understood from Table 1, where $R^{-1}$ is unstable if $\mathrm{ran}(R)\neq L_{2}^{q}$, since then some inputs in $L_{2}^{q}$ will map into $L_{2\mathrm{e}}^{p}\setminus L_{2}^{p}$. Additionally, it is not even clear if $R^{-1}$ exists at all. Note that all other operations in Theorem 2, these are multiplication with a real gain, addition with identity, parallel and series interconnection, all yield a stable system if the individual systems are stable.
In this section, we develop practical system analysis tools that solve these existence and stability problems that occur under the inversion operation. We first focus on the analysis of the troublesome part of a feedback system, i.e. where the inversion occurs. For the LFR in Eq. (11), this corresponds to the part $(\Phi^{-1}-G_{\mathrm{zw}})^{-1}$. We develop the theory for both the incremental and the non-incremental setting. Finally, we focus on systems in LFR form and provide a practical theorem that guarantees stability, well-posedness and incremental $L_{2}$-gain performance.
6.1 Incremental Stability Theorems
Consider the feedback interconnection in Fig. 1(a), where $H_{1}:L_{2}^{p}\to L_{2}^{q}$, $H_{2}:L_{2}^{q}\to L_{2}^{p}$, and the closed-loop operator reads $y=(H_{1}^{-1}+H_{2})^{-1}u$. As in [19], we abbreviate such operator interconnections as $(H_{1}^{-1}+H_{2})^{-1}=:[H_{1},H_{2}]$.
Definition 6 (Well-posedness).
Consider $H_{1}:L_{2\mathrm{e}}^{p}\to L_{2\mathrm{e}}^{q}$ and $H_{2}:L_{2\mathrm{e}}^{q}\to L_{2\mathrm{e}}^{p}$. We call the interconnection $[H_{1},H_{2}]$ well-posed if for all $u\in L_{2}^{p}$, there exist unique $e\in L_{2\mathrm{e}}^{p}$ and $y\in L_{2\mathrm{e}}^{q}$ such that $e=u-H_{2}y$ and $y=H_{1}e$.
Inspired by [28, Ch. 5], our well-posedness definition only assumes solutions to exist for $u\in L_{2}^{p}$, and not $u\in L_{2\mathrm{e}}^{p}$ as in [19]. Causality is not part of our well-posedness definition, as opposed to [29, 30]. This choice separates stability analysis on $L_{2}^{p}$ from causality, allowing for non-causal multipliers [31]. Therefore, Definition 6 imposes the minimal amount of structure upon the feedback system.
Let $n\geq\max\{p,q\}$, and embed $H_{1}$ and $H_{2}$ into $\mathscr{N}(L_{2}^{n})$ as $\tilde{H}_{1}=\iota_{n\leftarrow q}H_{1}\pi_{p\leftarrow n}$ and $\tilde{H}_{2}=\iota_{n\leftarrow p}H_{1}\pi_{q\leftarrow n}$, respectively. Note that these embeddings obey $\tilde{H}_{1}:\mathscr{U}_{p}^{n}\to\mathscr{U}_{q}^{n}$ and $\tilde{H}_{2}:\mathscr{U}_{q}^{n}\to\mathscr{U}_{p}^{n}$, where $\mathscr{U}_{p}^{n},\mathscr{U}_{q}^{n}$ are Banach spaces. This motivates the following statement of the incremental small gain theorem [28] on Banach spaces.
Lemma 5.
Let $H_{1}:\mathscr{U}\to\mathscr{Y}$, $H_{2}:\mathscr{Y}\to\mathscr{U}$, where $\mathscr{U},\mathscr{Y}$ are Banach spaces. If $\Gamma(H_{1})\Gamma(H_{2})<1$, then $[H_{1},H_{2}]:\mathscr{U}\to\mathscr{Y}$ is well-posed and incrementally stable.
The proof can be found in the Appendix.
The important feature of Lemma 5 is that it allows one to analyze a feedback interconnection in terms of $H_{1},H_{2}$, or their embeddings $\tilde{H}_{1},\tilde{H}_{2}$, on equal footing. To go beyond the small-gain theorem using SRGs, we need to phrase [19, Theorem 2] in the Banach space setting.
Lemma 6.
Let $H_{1}:\mathscr{U}\to\mathscr{Y}$ and $H_{2}:\mathscr{Y}\to\mathscr{U}$, where $\mathscr{U},\mathscr{Y}$ are Banach spaces, such that
  • $\Gamma(H_{1})<\infty$ and $\Gamma(H_{2})<\infty$,
  • $\exists\,\hat{\Gamma}>0$ such that $\Gamma([H_{1},\tau H_{2}])\leq\hat{\Gamma}$, for all $\tau\in[0,1]$.
Then, $[H_{1},\tau H_{2}]$ is well-posed for all $\tau\in[0,1]$ and incrementally stable with $\Gamma([H_{1},H_{2}])\leq\hat{\Gamma}$.
The proof can be found in the Appendix.
The calculus of SRGs, as developed in Section 5, provides us with a bound $\Gamma([\tilde{H}_{1},\tilde{H}_{2}])$, while $\mathrm{dom}([\tilde{H}_{1},\tilde{H}_{2}])=\mathrm{ran}(\tilde{H}_{1}^{-1}+\tilde{H}_{2})\subseteq\mathscr{U}_{p}^{n}$ is unknown. The practical use of Lemma 6 is to establish that $\mathrm{dom}([\tilde{H}_{1},\tilde{H}_{2}])=\mathscr{U}_{p}^{n}$. We are now in shape to state the core theorem for the analysis of incrementally stable systems using SRGs.
Theorem 4.
Consider the systems $H_{1}:L_{2}^{p}\to L_{2}^{q}$ and $H_{2}:L_{2}^{q}\to L_{2}^{p}$, where at least one of $\operatorname{SRG}(H_{1}),\operatorname{SRG}(H_{2})$ satisfies the chord property. If for all $\tau\in[0,1]$
$r_{\mathrm{min}}(\operatorname{SRG}(H_{1}))<\infty\text{ and }r_{\mathrm{min}}(\operatorname{SRG}(H_{2}))<\infty,$ (21a)
$\mathrm{dist}(\operatorname{SRG}(H_{1})^{-1},-\tau\operatorname{SRG}(H_{2}))\geq r>0,$ (21b)
then the interconnection $[H_{1},H_{2}]:L_{2}^{p}\to L_{2}^{q}$ is well-posed and incrementaly stable with $\Gamma([H_{1},H_{2}])\leq 1/r$. Moreover, if $H_{1}$ and $H_{2}$ are causal, then $[H_{1},H_{2}]$ is causal.
The proof can be found in the Appendix.
6.2 Non-Incremental Stability Theorems
Note that we assume throughout that $H_{1},H_{2}$ allow inputs in extended signal spaces, which is necessary for the definition of well-posedness, see Definition 6. This was not relevant in the incremental setting of Section 6.1, since well-posedness was obtained via the Banach fixed point theorem.
We state the non-incremental small gain theorem, and prove the non-incremental version of [19, Theorem 2].
Lemma 7.
Consider the systems $H_{1}:L_{2}^{p}\to L_{2}^{q}$ and $H_{2}:L_{2}^{q}\to L_{2}^{p}$. If $[H_{1},H_{2}]$ is well-posed and $\gamma(H_{1})\gamma(H_{2})<1$, then $[H_{1},H_{2}]:L_{2}^{p}\to L_{2}^{q}$.
Proof.
See [28, Theorem III.2.1]. ∎
Lemma 8.
Consider the systems $H_{1}:L_{2}^{p}\to L_{2}^{q}$ and $H_{2}:L_{2}^{q}\to L_{2}^{p}$. If for all $\tau\in[0,1]$ there exists a $\hat{\gamma}$ such that for all
  • $\gamma(H_{1})<\infty$ and $\gamma(H_{2})<\infty$,
  • $[H_{1},\tau H_{2}]$ is well-posed,
  • $\exists\hat{\gamma}>0$ such that $\gamma([H_{1},\tau H_{2}])\leq\hat{\gamma}$,
then $[H_{1},H_{2}]:L_{2}^{p}\to L_{2}^{q}$ is well-posed with $\gamma([H_{1},H_{2}])\leq\hat{\gamma}$.
The proof can be found in the Appendix.
Now we can state and prove the non-incremental analog of Theorem 4.
Theorem 5.
Consider the systems $H_{1}:L_{2}^{p}\to L_{2}^{q}$ and $H_{2}:L_{2}^{q}\to L_{2}^{p}$, where at least one of $\operatorname{SG}_{0}(H_{1}),\operatorname{SG}_{0}(H_{2})$ satisfies the chord property. If for all $\tau\in[0,1]$, the interconnection $[H_{1},\tau H_{2}]$ is well-posed and
$r_{\mathrm{min}}(\operatorname{SG}_{0}(H_{1}))<\infty\text{ and }r_{\mathrm{min}}(\operatorname{SG}_{0}(H_{2}))<\infty,$ (22a)
$\mathrm{dist}(\operatorname{SG}_{0}(H_{1})^{-1},-\tau\operatorname{SG}_{0}(H_{2}))\geq r>0,$ (22b)
then the interconnection $[H_{1},H_{2}]:L_{2}^{p}\to L_{2}^{q}$ is well-posed and non-incrementally stable with $\gamma([H_{1},H_{2}])\leq 1/r$.
The proof can be found in the Appendix.
6.3 Analysis of Systems in LFR Form
While the interconnection in Fig. 1(a) often captures the essential parts of analyzing a feedback system, which are stability and well-posedness, not all systems can be represented in this form. Instead, the LFR form described in Section 3 can describe a broader class of nonlinear systems.
From Eq. (11) we can see that the stability of the closed loop depends on the stability of $(\Phi^{-1}-G_{\mathrm{zw}})^{-1}$, i.e.
$[\Phi,-G_{\mathrm{zw}}]:\mathrm{dom}([\Phi,-G_{\mathrm{zw}}])\subseteq L_{2}^{n_{\mathrm{z}}}\to L_{2}^{n_{\mathrm{w}}},$ (23)
which is precisely what we can analyze using Theorem 4. We call such an LFR system well-posed if Eq. (23) is well-posed.
The incremental gain $\Gamma(R)$ of an LFR system is obtained by replacing the operators in Eq. (11) with their SRG, and computing the radius of the resulting set.
Theorem 6.
Consider the system $R:\mathrm{dom}(R)\subseteq L_{2}^{p}\to L_{2}^{q}$ given by the LFR in Eq. (11), where $G\in RH_{\infty}^{(q+n_{\mathrm{z}})\times(p+n_{\mathrm{w}})}$ and $\Phi:L_{2}^{n_{\mathrm{z}}}\to L_{2}^{n_{\mathrm{w}}}$ satisfy $\Gamma(G)<\infty$ and $\Gamma(\Phi)<\infty$. If there exists a $\hat{\Gamma}<\infty$ such that $\forall\,\tau\in[0,1]$
$r_{\mathrm{min}}(\mathscr{G}_{R})\leq\hat{\Gamma},$ (24a)
$\mathscr{G}_{R}:=\overline{\operatorname{SRG}(G_{\mathrm{yu}})+\operatorname{SRG}(G_{\mathrm{yw}})}\\ \overline{\times(\operatorname{SRG}(\Phi)^{-1}-\tau\operatorname{SRG}(G_{\mathrm{zw}}))^{-1}\operatorname{SRG}(G_{\mathrm{zu}})},$ (24b)
where the overline indicates the bound obtained in Eq. (20), then $R:L_{2}^{p}\to L_{2}^{q}$ is well-posed and incrementally stable with $\Gamma(R)\leq\hat{\Gamma}$. If $\Phi$ is causal, then so is $R$.
The proof can be found in the Appendix.
Remark 3.
Theorem 6 can be restated in the non-incremental setting if $\Phi(0)=0$ and assuming that $[\Phi,-\tau G_{\mathrm{zw}}]$ is well-posed for all $\tau\in[0,1]$. Furthermore, in this case, causality of $G$ and $\Phi$ no longer implies causality of $R$.
Remark 4.
It is not necessary to use the improved chord/arc completions. It is sufficient when the chord (arc) property is satisfied for each sum (product) in Eq. (24). However, this may lead to a larger value of $\hat{\Gamma}$, i.e. more conservatism.
Note that $\Phi$ in Theorem 6 may be any nonlinear operator with a finite SRG bound, not just a diagonal static nonlinear block. Examples of dynamic nonlinearities are time-varying static nonlinearities such as $x\mapsto\sin(t)\sin(x)$, and hybrid systems. For the latter, SG bounds can be found in [9, 32].
In this section, we have developed stability theorems for the feedback interconnections in Fig. 1. We note, however, that the methods developed in this paper can be used to analyze any interconnection of MIMO systems by using the modular interconnection rules from Section 5, and the tools from Section 6 to analyze the feedback loops.
7 Computing the MIMO SRG of Common Operators
Now, all theoretical tools to analyze the LFR system in Eq. (11) are in place. To perform the analysis of a given system, one requires expressions for the SRG of the operators $G$ and $\Phi$ in Eqs. (9) and (10), respectively.
For this purpose, we develop formulas to compute an SRG bound of two common MIMO operators: non-square stable LTI operators ant the class of nonlinear operators, which are diagonal and static. These results provide the tools to compute the MIMO SRGs of all operators in Eq. (11), but they can also be used in different interconnection architectures.
7.1 Computing the MIMO SRG of LTI Operators
The first ingredient of a system in LFR form, see Eq. (11), is an LTI operator. We will show how to calculate a bound for the SRG of a MIMO LTI operator, as defined in Definition 3. Let $G\in RH_{\infty}^{q\times p}$ be the transfer function $G(s)$ that corresponds to a stable causal LTI operator $G:L_{2}^{p}\to L_{2}^{q}$, i.e. $G(s)$ is proper and all poles $p$ obey $\mathrm{Re}(p)<0$.
7.1.1 Upper bounding the SRG
Recall that the $H_{\infty}$ norm has the property that
$\left\lVert G\right\rVert_{\infty}=\sup_{u\in L_{2}^{p}}\frac{\left\lVert Gu\right\rVert}{\left\lVert u\right\rVert}=\sup_{\omega\in\mathbb{R}}\overline{\sigma}(G(j\omega))=:\overline{\sigma}(G),$
hence $\operatorname{SRG}(G)\subseteq D_{\overline{\sigma}(G)}(0)$. Now define $n=\max\{p,q\}$ and
$G_{\alpha}=\begin{pmatrix}G\\ 0_{(n-q)\times p}\end{pmatrix}-\begin{pmatrix}\alpha I\\ 0_{(n-p)\times p}\end{pmatrix},$ (25)
where $I\in\mathbb{R}^{p\times p}$ is the identity, $\alpha\in\mathbb{R}$ and $G_{\alpha}:L_{2}^{p}\to L_{2}^{n}$. Observe that
$G_{\alpha}\pi=\iota G\pi-\begin{pmatrix}\alpha I&0\\ 0&0\end{pmatrix}=\iota G\pi-\alpha I_{\mathscr{U}_{p}^{n}},$
where $\pi=\pi_{p\leftarrow n}$ and $\iota=\iota_{n\leftarrow p}$. Since the last $n-p$ inputs are zero in $\mathscr{U}_{p}^{n}$, we know from the SRG definition that
$\operatorname{SRG}_{\mathscr{U}_{p}^{n}}(G_{\alpha}\pi_{p\leftarrow n})\subseteq D_{\upsilon_{\alpha}}(0),\quad\upsilon_{\alpha}:=\overline{\sigma}(G_{\alpha}).$ (26)
Using Theorem 2.a. and 2.b. on Eq. (26) we can conclude
$\operatorname{SRG}_{\mathscr{U}_{p}^{n}}(\iota G\pi)\subseteq D_{\upsilon_{\alpha}}(\alpha),$ (27)
hence $\operatorname{SRG}(G)\subseteq D_{\upsilon_{\alpha}}(\alpha)$ by Definition 3. For some set $\Upsilon\subseteq\mathbb{R}$, we have the SRG bound
$\operatorname{SRG}(G)\subseteq\bigcap_{\alpha\in\Upsilon}D_{\upsilon_{\alpha}}(\alpha).$ (28)
7.1.2 Lower bounding the SRG
Let $\underline{\sigma}(G):=\inf_{\omega\in\mathbb{R}}\underline{\sigma}(G(j\omega))$ be the smallest singular value of $G(s)$ on the imaginary axis. By Parseval’s theorem, one has $\left\lVert Gu\right\rVert\geq\underline{\sigma}(G)\left\lVert u\right\rVert$ for all $u\in L_{2}^{p}$, and therefore
$\operatorname{SRG}(G)\subseteq\mathbb{C}_{\infty}\setminus D_{\underline{\sigma}(G)}(0).$
Denote the lower bound $\ell_{\alpha}:=\underline{\sigma}(G_{\alpha})$, where $G_{\alpha}$ is taken from Eq. (25). By the same reasoning used to derive the upper bound in Eq. (28), we have the following SRG lower bound
$\operatorname{SRG}(G)\subseteq\mathbb{C}_{\infty}\setminus\bigcup_{\alpha\in\Lambda}D_{\ell_{\alpha}}(\alpha)$ (29)
where $\Lambda\subseteq\mathbb{R}$.
7.1.3 Algorithm to bound the SRG of LTI operators
We can now state one of our main results, which is an SRG bound for MIMO LTI operators.
Theorem 7.
Let $G\in RH_{\infty}^{q\times p}$ and $\Upsilon,\Lambda\subseteq\mathbb{R}$, then
$\operatorname{SRG}(G)\subseteq\Bigl{(}\bigcap_{\alpha\in\Upsilon}D_{\upsilon_{\alpha}}(\alpha)\Bigl{)}\setminus\Bigl{(}\bigcup_{\alpha\in\Lambda}D_{\ell_{\alpha}}(\alpha)\Bigl{)}=:\mathscr{G}^{G}_{\Upsilon,\Lambda}.$ (30)
Proof.
The result follows directly from Eqs. (28) and (29). ∎
The result in Theorem 7 has an intuitive interpretation in terms of disks in the complex plane. For each $\alpha\in\Upsilon$, using $\overline{\sigma}(G_{\alpha})$, we compute the disk $D_{\upsilon_{\alpha}}(\alpha)$ centered at $\alpha$ that contains $\operatorname{SRG}(G)$. Similarly, for each $\alpha\in\Lambda$, using $\underline{\sigma}(G_{\alpha})$ we compute the $D_{\ell_{\alpha}}(\alpha)$ centered at $\alpha$ which does not contain $\operatorname{SRG}(G)$. In Eq. (30), we then intersect all disks that contain $\operatorname{SRG}(G)$, and remove all disks that do not contain $\operatorname{SRG}(G)$.
If $|\Upsilon|=n_{\upsilon},|\Lambda|=n_{\ell}$, then Eq. (30) amounts to $n_{\upsilon}+n_{\ell}$ amount of $H_{\infty}$ norm computations. To represent $\mathscr{G}^{G}_{\Upsilon,\Lambda}$, which consists of the union/intersection of $n_{\upsilon}+n_{\ell}$ circles, each requiring a radius and center on $\mathbb{R}$ to be uniquely represented, one needs to store only $2(n_{\upsilon}+n_{\ell})$ real numbers.
By abuse of notation, we will use $\operatorname{SRG}(G)=\mathscr{G}^{G}_{\Upsilon,\Lambda}$. In practice, we often take $\Upsilon=\Lambda$, i.e. the circles for maximum and minimum gain are computed at the same base points for computational efficiency. An algorithm for computing the bound in Eq. (30) is available at github.com/Krebbekx/SrgTools.jl.
An algorithm for bounding the SRG of an LTI system, based on shifting circles on the real axis, has been proposed before [32]. However, this method works only for SISO or square MIMO systems since they are based on IQCs. By the same token, it is not possible to compute the SRG on a restricted set of inputs of certain frequency using [32], whereas Theorem 7 allows this by simply restricting the frequency interval over which the largest/smallest singular value is computed. The latter is useful to study the frequency domain performance using SRGs [7] in practice. From [32], we can conclude that Eq. (30) becomes an equality for square systems if $\Upsilon=\Lambda=\mathbb{R}$.
Example 1.
Consider the transfer functions
$G_{1}(s)$ $=\begin{pmatrix}\frac{s}{s+1}&\frac{s^{2}}{s^{2}+s+1}&\frac{1}{2s+1}\end{pmatrix},$
$G_{2}(s)$ $=\begin{pmatrix}\frac{s^{2}}{s^{2}+s+1}&\frac{1}{2s+1}\\ \frac{s+1}{(s+3)(s^{2}+s+1)}&\frac{s+3}{s+1}\\ \frac{s^{2}-1}{(s+3)(s+2)}&\frac{s}{s+2}\end{pmatrix}.$
Their SRGs, computed with Theorem 7 are shown in Fig. 2.
(a) $G_{1}(s)$.
(b) $G_{2}(s)$.
Figure 2: SRGs of the MIMO transfer functions in Example 1.
Since the identity operator acts differently on each row/column of Eq. (25), it is clear that the direction of inputs and outputs influences the bound Eq. (30). This shows that the MIMO LTI SRG is not agnostic to input/output directions. This means one can perform SRG calculations for all possible permutations of inputs/outputs in the system and choose the least conservative result. Therefore, input/output directions can be seen as a degree of freedom in the analysis with SRGs.
Note that Theorem 7 also provides an effective algorithm to compute the SRG of a matrix. In that case, the search over all frequencies for the largest/smallest singular value can be skipped as the matrix can be viewed as a transfer matrix with no frequency dependence. This approach can be used to generalize [13, 14, 15] to the case of non-square LTI systems.
7.2 Computing the MIMO SRG of Diagonal Static Nonlinear Operators
The second component of a system in LFR form, see Eq. (11), is a nonlinear operator. We consider a square operator $\Phi:L_{2}^{d}\to L_{2}^{d}$ that is defined by the diagonal nonlinear map
$\mathbb{R}^{d}\ni(x_{1},\dots,x_{d})^{\top}\mapsto(\phi_{1}(x_{1}),\dots,\phi_{d}(x_{d}))^{\top}\in\mathbb{R}^{d},$ (31)
where $\phi_{i}:\mathbb{R}\to\mathbb{R}$ is an incrementally sector bounded static nonlinear map defined by
$\mu_{i}|x-y|^{2}\leq(x-y)(\phi_{i}(x)-\phi_{i}(y))\leq\lambda_{i}|x-y|^{2},\;\forall x,y\in\mathbb{R},$
where $\mu_{i},\lambda_{i}\in\mathbb{R}$, which is denoted as $\partial\phi_{i}\in[\mu_{i},\lambda_{i}]$.
Lemma 9.
Let $\Phi:L_{2}^{d}\to L_{2}^{d}$ be defined by Eq. (31), where $\partial\phi_{i}\in[\mu_{i},\lambda_{i}]$ and $\mu=\min_{i}\mu_{i}$ and $\lambda=\max_{i}\lambda_{i}$, then
$\operatorname{SRG}(\Phi)\subseteq D_{r}(c)=D_{[\mu,\lambda]},$ (32)
with center $c=\frac{\lambda+\mu}{2}$ and radius $r=\frac{\lambda-\mu}{2}$.
The proof can be found in the Appendix.
A similar result can be obtained for non-incremental sector bounded operators, i.e. if the operators $\phi_{i}$ in Eq. (31) obey
$\mu_{i}|x|^{2}\leq x\phi_{i}(x)\leq\lambda_{i}|x|^{2},\;\forall x\in\mathbb{R}$
where $\mu_{i},\lambda_{i}\in\mathbb{R}$, which is denoted as $\phi_{i}\in[\mu_{i},\lambda_{i}]$.
Lemma 10.
Let $\Phi:L_{2}^{d}\to L_{2}^{d}$ be defined by Eq. (31), where $\phi_{i}\in[\mu_{i},\lambda_{i}]$ and $\mu=\min_{i}\mu_{i}$ and $\lambda=\max_{i}\lambda_{i}$, then
$\operatorname{SG}_{0}(\Phi)\subseteq D_{r}(c)=D_{[\mu,\lambda]},$ (33)
with center $c=\frac{\lambda+\mu}{2}$ and radius $r=\frac{\lambda-\mu}{2}$.
Proof.
The result is obtained by fixing $y=0$ in Lemma 9. ∎
Remark 5.
Lemma 9 only takes the smallest $\mu_{i}$ and largest $\lambda_{i}$ into account. Therefore, one could reduce the conservatism of SRG calculations by applying loop transformations such that $\mu=\mu_{i},\lambda=\lambda_{i}$ for all $i=1,\dots,d$.
8 Examples
With the following examples we demonstrate how the system analysis results from Section 6 can be used to analyze the stability and $L_{2}$-gain performance of example systems. Throughout, we use the methods from Section 7 to obtain SRG bounds for the operators that are involved.
For each example, the Julia code used to generate all figures and other results is freely available at github.com/Krebbekx/SrgTools.jl.
8.1 SISO System with Multiple Nonlinearities
$\Sigma$$K(s)$$\phi_{1}$$\Sigma$$P(s)$$\phi_{2}$$u$$e$$u$$\hat{u}$$u^{\prime}$$y$----
Figure 3: Block diagram of a controlled Lur’e plant with saturation.
(a) $\operatorname{SRG}(G_{\mathrm{yw}})$
(b) $\operatorname{SRG}(G_{\mathrm{zu}})$
(c) $\operatorname{SRG}(G_{\mathrm{yu}})$
(d) $\mathscr{G}_{R}$
Figure 4: SRG analysis of the example in Section 8.1.
(a) SRGs of $\Phi$ and $G_{\mathrm{zw}}$.
(b) SRGs of $\Phi^{a}$ and $G_{\mathrm{zw}}^{a}$.
Figure 5: SRG analysis of the example in Section 8.1.
Consider the system in Fig. 3, where $P,K,\phi_{1},\phi_{2}:L_{2\mathrm{e}}\to L_{2\mathrm{e}}$ are causal SISO systems defined as
$\begin{split}K(s)=\frac{1}{s+1},\quad P(s)=\frac{3}{(s-2)(s/10+1)},\\ \phi_{1}(s)=\begin{cases}x\text{ if }|x|\leq 1,\\ x/|x|\text{ else},\end{cases}\phi_{2}(x)=\begin{cases}x\text{ if }|x|\leq 1,\\ 2x-x/|x|\text{ else}.\end{cases}\end{split}$
which is also studied as an example in [6]. Define the loop transformations $\varphi_{1}:=\phi_{1}-\kappa_{1}$ and $\varphi_{2}:=\phi_{2}-\kappa_{2}$, where $\kappa_{1},\kappa_{2}\in\mathbb{R}$. To write this system in LFR form $y=Ru$ according to Eq. (11), the nonlinearity becomes $\Phi:L_{2\mathrm{e}}^{2}\to L_{2\mathrm{e}}^{2}$ defined by $(x,y)\mapsto(\varphi_{1}(x),\varphi_{2}(y))$. The LTI part in Eq. (9) is given by
$G_{\mathrm{zw}}$ $=\begin{pmatrix}-S\tilde{P}K&-S\tilde{P}K\\ S\tilde{P}&S\tilde{P}\end{pmatrix},\quad$ $G_{\mathrm{zu}}=\begin{pmatrix}SK\\ SL\end{pmatrix},$ (34)
$G_{\mathrm{yw}}$ $=\begin{pmatrix}S\tilde{P}&S\tilde{P}\end{pmatrix},$ $G_{\mathrm{yu}}=SL,$
where $\tilde{P}=\frac{P}{1+\kappa_{2}P},L=\kappa_{1}\tilde{P}K$ and $S=\frac{1}{1+L}$. Note that $\partial\phi_{1}\in[0,1]$ and $\partial\phi_{2}\in[1,2]$, therefore $\partial\varphi_{1}\in[-\kappa_{1},1-\kappa_{1}]$ and $\partial\varphi_{2}\in[1-\kappa_{2},2-\kappa_{2}]$. Then, according to Lemma 9, we have the SRG bound
$\operatorname{SRG}(\Phi)\subseteq D_{[\min\{-\kappa_{1},1-\kappa_{2}\},\max\{1-\kappa_{1},2-\kappa_{2}\}]}.$
Before we can apply Theorem 6, we must make sure that $G$, defined by Eq. (9) and the transfer functions in Eq. (34), is stable. This is achieved by picking loop transformation variables $\kappa_{1},\kappa_{2}$ such that $G$ is stable.
We fix $\kappa_{1}=2,\kappa_{2}=3$, for which all poles $p$ of $G$ satisfy $\mathrm{Re}(p)<0$. As an alternative choice, denoted by a superscript $\mathrm{a}$, we consider is $\kappa_{1}^{\mathrm{a}}=0.5,\kappa_{2}^{\mathrm{a}}=1.5$, for which $G^{\mathrm{a}}$ is stable as well. Note that these choices result in $\partial\varphi_{1},\partial\varphi_{2}\in[-2,-1]$ and $\partial\varphi_{1}^{\mathrm{a}},\partial\varphi_{2}^{\mathrm{a}}\in[-0.5,0.5]$ for the loop-transformed nonlinearities. We will now apply Theorem 6 for both choices of the loop transformation variables.
The first step is to compute $\operatorname{SRG}(G_{\mathrm{zw}})$ using Theorem 7. Since $\partial\varphi_{1},\partial\varphi_{2}\in[-2,-1]$, we can conclude that $\operatorname{SRG}(\Phi)\subseteq D_{[-2,-1]}$. The stability of $R$ depends on $[\Phi,-G_{\mathrm{zw}}]$, is equivalent to the requirement that, for all $\tau\in[0,1]$, $\operatorname{SRG}(\Phi)^{-1}$ and $\tau\operatorname{SRG}(G_{\mathrm{zw}})$ do not overlap. These graphs are plotted in Fig. 5(a) and are indeed separated for all $\tau\in[0,1]$. Analogously, for the alternative choice of loop transformation variables, the graphs $\operatorname{SRG}(\Phi^{\mathrm{a}})^{-1}$ and $\tau\operatorname{SRG}(G_{\mathrm{zw}}^{\mathrm{a}})$ are visualized in Fig. 5(b). From Fig. 5 it is clear that the smallest separation is attained at $\tau=1$.
The SRGs of $G_{\mathrm{yw}},G_{\mathrm{zu}}$ and $G_{\mathrm{yu}}$ are also computed using Theorem 7, and visualized in Figs. 4(a), 4(b) and 4(c), respectively. The final step is to evaluate $\mathscr{G}_{R}$ in Eq. (24) using the SRG interconnection rules in Theorem 2. In each step, we applied the improved chord and arc completions from Lemma 4, yielding the set $\mathscr{G}_{R}$ in Fig. 4(d). Since $r_{\mathrm{min}}(\mathscr{G}_{R})\leq 2.33$, we can conclude that $R:L_{2\mathrm{e}}\to L_{2\mathrm{e}}$ is a well-posed and causal system which satisfies $\Gamma(R)\leq 2.33$.
A similar computation for the alternative loop transformation variables yields $r_{\mathrm{min}}(\mathscr{G}_{R}^{\mathrm{a}})\leq 6.13$. This shows that the outcome of Theorem 6 can be optimized over all choices of loop transformations that stabilize $G$.
This example is also studied in [6], where a bound $\Gamma(R)\leq 4.81$ is obtained by using SISO SRG tools only. Hence, we see that our MIMO approach yields a tighter incremental $L_{2}$-gain bound. Moreover, we obtain causality and well-posedness of $R$ via Theorem 6, which are both properties that had to be assumed in [6].
8.2 Two Mass-Spring-Damper System
We consider a system of two masses $m_{1}$ and $m_{2}$ that are connected to the solid ground and each other with a linear spring and damper, and a nonlinear spring, as depicted in Fig. 6. We take external forces $u_{1},u_{2}$ on the masses $m_{1},m_{2}$ as inputs and as outputs the positions $x_{1},x_{2}$ of the masses. The system is governed by
$m_{1}\ddot{x}_{1}$ $=u_{1}-k_{1}x_{1}-d_{1}\dot{x}_{1}-k_{12}(x_{1}-x_{2})-d_{12}(\dot{x}_{1}-\dot{x}_{2})$ (35)
$+\phi_{1}(x_{1})+\phi_{12}(x_{1}-x_{2}),$
$m_{2}\ddot{x}_{2}$ $=u_{2}-k_{2}x_{2}-d_{2}\dot{x}_{2}+k_{12}(x_{1}-x_{2})+d_{12}(\dot{x}_{1}-\dot{x}_{2})$
$+\phi_{2}(x_{2})-\phi_{12}(x_{1}-x_{2}),$
where $k_{1},k_{2}$ and $d_{1},d_{2}$ are the linear spring and damper coefficients, respectively. We choose parameters $m_{1}=0.5,m_{2}=3,k_{1}=1,k_{2}=2,d_{1}=0.3,d_{2}=1,d_{12}=1,k_{12}=0.5$ and nonlinear springs $\phi_{1}(x)=\phi_{2}(x)=\phi(x):=-\tanh(x)$ and $\phi_{12}(x)=2\tanh(x)-x$. Hence, $\phi$ is a saturating spring and $\phi_{12}$ is a negative spring for small deflection, and a regular spring for large deflection.
Figure 6: The nonlinear mass-spring-damper setup.
To write Eq. (35) in LFR form Eq. (11), we take the nonlinear function $\Phi(x,y,z)=(\phi(x),\phi(y),\phi_{12}(z))^{\top}$, and signals $u=(u_{1},u_{2})^{\top},w=(\phi(x_{1}),\phi(x_{2}),\phi_{12}(x_{1}-x_{2}))^{\top},y=(x_{1},x_{2})^{\top}$ and $z=(x_{1},x_{2},x_{1}-x_{2})^{\top}$. The transfer function $G$ in Eq. (9) is obtained from the state-space representation
[…]
Since $\partial\phi\in[-1,0]$ and $\partial\phi_{12}\in[-1,1]$, which are not the same intervals, we want to shift and scale $\phi$ and $\phi_{12}$ such that they satisfy the same sector bound in order to tighten the bound in Lemma 9. To that end, we define $\varphi(x)=\phi(x)+\frac{1}{2}x$ and $\varphi_{12}(x)=\frac{1}{2}\phi_{12}(x)$ such that $\partial\varphi\in[-\frac{1}{2},\frac{1}{2}]$ and $\partial\varphi_{12}\in[-\frac{1}{2},\frac{1}{2}]$. To accommodate for this loop transformation in the LFR, we define $\tilde{G}(s)$ by $\tilde{k}_{1}=k_{1}+\frac{1}{2},\tilde{k}_{2}=k_{2}+\frac{1}{2}$ and $\tilde{G}_{\mathrm{x}w_{3}}=2G_{\mathrm{x}w_{3}}$ where $\mathrm{x}\in\{y_{1},y_{2},z_{1},z_{2},z_{3}\}$, and all other values are the same as for $G(s)$. The nonlinear function in Eq. (11) becomes $\tilde{\Phi}(x,y,z)=(\varphi(x),\varphi(y),\varphi_{12}(z))^{\top}$, hence $\operatorname{SRG}(\tilde{\Phi})\subseteq D_{[-\frac{1}{2},\frac{1}{2}]}$ by Lemma 9.
(a) SRGs of $\tilde{\Phi}$ and $\tilde{G}_{\mathrm{zw}}$.
(b) SRG bound $\mathscr{G}_{R}$ from Eq. (24).
Figure 7: SRG analysis of the example in Section 8.2.
To apply Theorem 6 we must first apply the loop transformation to obtain $\tilde{G}(s)$ and check if $\tilde{G}(s)\in RH_{\infty}^{5\times 5}$, which is the case. Second, we check stability of $[\tilde{\Phi},-\tilde{G}_{\mathrm{zw}}]$ by plotting their SRGs, see Fig 7(a), where is clear that $\tau\operatorname{SRG}(\tilde{G}_{\mathrm{zw}})$ and $\operatorname{SRG}(\tilde{\Phi})^{-1}$ do not touch for all $\tau\in[0,1]$. Finally, we compute the SRG bound in Eq. (24) which yields $\hat{\Gamma}(R)=12.09$, see Fig. 7(b). From Theorem 6, we can conclude that the system is causal and well-posed with incremental $L_{2}$-gain bound $\Gamma(R)\leq 12.09$. The large gain bound can be understood by the fact that the spring between $m_{1}$ and $m_{2}$ is negative, i.e. active, for small deflections. Therefore, for small inputs $u_{1},u_{2}$, relatively large outputs $x_{1},x_{2}$ can be expected. Also, we did not yet optimize the bound $\hat{\Gamma}$ over all possible loop transformations.
Note that in this example, we have used a loop transformation that consists of both a shift in elements of $\Phi$, and a scaling between the output of $\Phi$ and input of $G(s)$.
8.3 Comparison with IQC Based SRG Results
Figure 8: Feedback diagram of the example in Section 8.3 (image taken from [18]).
We will now treat an example where non-incremental stability is analyzed. Consider the feedback diagram in Fig. 8 where
$P(s)$ $=\begin{pmatrix}\frac{0.1}{s+1}&\frac{1}{s^{3}+5s^{2}+2s+1}\\ \frac{0.1}{s^{3}+5s^{2}+2s+1}&\frac{0.2}{s+5}\end{pmatrix},$
$H_{2}(s)$ $=\begin{pmatrix}\frac{1.7}{s^{2}+2s+1}&0\\ 0&\frac{1.7}{s^{2}+3s+3}\end{pmatrix},\quad\operatorname{SG}_{0}(\Phi)\subseteq D_{\sqrt{0.1}}(0),$
i.e. the nonlinearity $\Phi$ is any operator that satisfies $\gamma(\Phi)\leq\sqrt{0.1}$. This system is studied in [18], where a bound for the SG of $H_{1}:=[P,-\Phi]$ is obtained using an IQC-based method.
The SRG of $P$, obtained using Theorem 7, is plotted in Fig. 9(a). To analyze $H_{1}=[P,-\Phi]$ using Theorem 5, we plot $SRG(P)^{-1}$ and $\operatorname{SG}_{0}(\Phi)$ in Fig. 9(b). Since $\operatorname{SG}_{0}(\Phi)$ is a disk and hence satisfies the chord property, we can compute the bound $(\operatorname{SRG}(P)^{-1}-\operatorname{SG}_{0}(\Phi))^{-1}$ for $\operatorname{SG}_{0}(H_{1})$ using Theorem 3, which is shown in Fig. 9(c). We note that the SG bound in Fig. 9(c) is identical to the result obtained in [18], which uses an IQC-based analysis approach.
(a) $\operatorname{SRG}(P)$
(b) $\operatorname{SRG}(P)^{-1}$ and $\operatorname{SG}_{0}(\Phi)$
(c) $\operatorname{SG}_{0}(H_{1})$
(d) $\operatorname{SG}_{0}(T)$
Figure 9: SG analysis of example in Section 8.3.
We can also compute an SG bound for $T:=[H_{1},H_{2}]$ by noting that $[H_{1},H_{2}]=[G,-\Phi]$ where $G=[P,H_{2}]$. The SRG of $G$ is computed with Theorem 7, and the stability analysis is done using Theorem 5, entirely analogous to the analysis of $H_{1}$. The resulting SG bound for $T$ is shown in Fig. 9(d), from which we conclude that $\gamma(T)\leq r_{\mathrm{min}}(\operatorname{SG}_{0}(T))\leq 1.79$. In [18], an explicit value of the gain bound is not computed, but from their results it can be deduced that they obtain $\gamma(T)\leq\sqrt{16.38}\approx 4.05$ in their approach.
Since we work non-incrementally, we must assume that the systems $H_{1}$ and $T$ are well-posed in the sense of Definition 6.
9 Conclusion
In this paper we have developed the theoretical backbone of SRG analysis for nonlinear multivariable feedback systems that are interconnections of (possibly non-square) MIMO systems. We started by constructing an embedding for general MIMO operators into a space of square maps, while restricting the inputs to a relevant subspace. Next, we showed under which conditions MIMO operators can be interconnected, and provide interconnection rules for both the incremental and non-incremental case. We then restricted ourselves to MIMO systems in LFR form and provided practical stability and $L_{2}$-gain performance results. For operators in the LFR form, which are stable LTI operators that may be non-square, and diagonal static nonlinear operators, we provided explicit formulas for their MIMO SRG. Finally, we demonstrated our result on three examples; a SISO system with multiple nonlinearities, a nonlinear MIMO mass-spring-damper system, and a MIMO feedback system from [18]. The advantage of our framework is that if the input/output dimensions are compatible, the user can perform computations with the MIMO SRG without carrying out the embedding explicitly. Even though this paper focused on systems in LFR form, core results of our framework can be applied to any interconnection of MIMO systems.
Throughout the paper, we mentioned that there are certain degrees of freedom in MIMO SRG calculations. These are loop transformations that shift, scale and possibly permute the inputs and outputs of MIMO systems. In our first example we show that a different $L_{2}$-gain bound is obtained from two different choices of shifts in the loop transformation. In the second example we use shifts and a scaling to make sure that all nonlinearities lie in the same sector. A promising avenue of further work is optimizing the resulting $L_{2}$-gain upper bound over all possible loop transformations. The first step in this direction is explored in [18] for diagonal scalings in square systems. A second direction of interest is to compare the results of the MIMO SRG framework with IQC analysis, and possibly combine them to minimize the conservatism of SRG analysis.
[…]
Appendix
Proof of Lemma 1.
[…]
Proof of Lemma 2.
By Lemma 1, we know that $\iota_{q\leftarrow p}:L_{2}^{p}\to\mathscr{U}_{p}^{q}$ is an isomorphism. Since for all $u\in L_{2}^{p}$ and $\tilde{u}\in\mathscr{U}_{p}^{q}$ it holds that $\pi_{p\leftarrow q}\iota_{q\leftarrow p}u=u$ and $\iota_{q\leftarrow p}\pi_{p\leftarrow q}\tilde{u}=\tilde{u}$, we have shown that $\iota_{q\leftarrow p}^{-1}=\pi_{p\leftarrow q}$, hence $\pi_{p\leftarrow q}^{-1}=\iota_{q\leftarrow p}$. ∎
Proof of Lemma 3.
Denote $\iota:=\iota_{\tilde{n}\leftarrow n}$ and $\pi:=\pi_{n\leftarrow\tilde{n}}$ for brevity. By Lemma 1 and 2, we know that $\iota$ is an isometric isomorphism with isometric isomorphic inverse $\pi$. Therefore, $u_{1},u_{2}\in\mathscr{U}\iff\tilde{u}_{1},\tilde{u}_{2}\in\iota(\mathscr{U})$ where $\tilde{u}_{1}=\iota(u),\tilde{u}_{2}=\iota(u_{2})$ satisfying $\left\lVert u_{1}-u_{2}\right\rVert=\left\lVert\tilde{u}_{1}-\tilde{u}_{2}\right\rVert$ and
$\left\lVert\iota R\pi\tilde{u}_{1}-\iota R\pi\tilde{u}_{2}\right\rVert=\left\lVert\iota Ru_{1}-\iota R\pi u_{2}\right\rVert=\left\lVert Ru_{1}-R\pi u_{2}\right\rVert,\\ \left\langle\iota R\pi\tilde{u}_{1}-\iota R\pi\tilde{u}_{2},\tilde{u}_{1}-\tilde{u}_{2}\right\rangle=\left\langle\iota Ru_{1}-\iota Ru_{2},\tilde{u}_{1}-\tilde{u}_{2}\right\rangle\\ =\left\langle Ru_{1}-Ru_{2},u_{1}-u_{2}\right\rangle.$
This shows that $z_{R}\subseteq\operatorname{SRG}_{\mathscr{U}}(R)\iff z_{R}\subseteq\operatorname{SRG}_{\iota(\mathscr{U})}(\iota R\pi)$, where $z=z_{R}(u_{1},u_{2})=z_{R}(\tilde{u}_{1},\tilde{u}_{2})$, proving $\operatorname{SRG}_{\mathscr{U}}(R)=\operatorname{SRG}_{\iota(\mathscr{U})}(\iota R\pi)$. ∎
Proof of Theorem 2.
The proof extends [4], where we now take the details of the domain and range into account. We prove each point separately.
  1. a.
    As $\mathscr{U}$ is a linear subspace, we have that for any $u\in\mathscr{U}\implies\alpha u\in\mathscr{U}$. By Eq. (7) we have $\angle(\alpha u,y)=\angle(u,\alpha y)=\angle(u,y)$. We have $\operatorname{SRG}_{\mathscr{U}}(\alpha R)=\alpha\operatorname{SRG}_{\mathscr{U}}(R)$ per definition and since $(1/a)u\in\mathscr{U}$, we also have $\operatorname{SRG}_{\mathscr{U}}(R\alpha)=\alpha\operatorname{SRG}_{\mathscr{U}}(R)$.
  2. b.
    From [4] we use $\mathrm{Re}\,z_{R}(u_{1},u_{2})=\frac{\left\langle Ru_{1}-Ru_{2},u_{1}-u_{2}\right\rangle}{\left\lVert u_{1}-u_{2}\right\rVert^{2}}$ and $\mathrm{Im}\,z_{R}(u_{1},u_{2})=\pm\frac{\left\lVert\pi_{(u_{1}-u_{2})^{\perp}}(Ru_{1}-Ru_{2})\right\rVert}{\left\lVert u_{1}-u_{2}\right\rVert}$, where $\pi_{x^{\perp}}$ is the projection on the subspace orthogonal to $x$. Since for all $u_{1},u_{2}\in\mathscr{U}$ it holds that $\left\langle(I_{\mathscr{U}}+R)u_{1}-(I_{\mathscr{U}}+R)u_{2},u_{1}-u_{2}\right\rangle=\left\lVert u_{1}-u_{2}\right\rVert^{2}+\left\langle Ru_{1}-Ru_{2},u_{1}-u_{2}\right\rangle$ and $\pi_{(u_{1}-u_{2})^{\perp}}((I_{\mathscr{U}}+R)u_{1}-(I_{\mathscr{U}}+R)u_{2})=\pi_{(u_{1}-u_{2})^{\perp}}(Ru_{1}-Ru_{2})$, from which it follows that $\operatorname{SRG}_{\mathscr{U}}(I_{\mathscr{U}}+R)=1+\operatorname{SRG}_{\mathscr{U}}(R)$.
  3. c.
    Per definition of the Möbius inverse $re^{j\phi}\mapsto(1/r)e^{j\phi}$ one has $(\operatorname{SRG}_{\mathscr{U}}(R))^{-1}\setminus\{0,\infty\}=\{\frac{\left\lVert u_{1}-u_{2}\right\rVert}{\left\lVert y_{1}-y_{2}\right\rVert}e^{\pm j\angle(u_{1}-u_{2},y_{1}-y_{2})}\mid(u_{1},y_{1}),(u_{2},y_{2})\in R,u_{1}\neq u_{2},y_{1}\neq y_{2},u_{1},u_{2}\in\mathscr{U}\}=\{\frac{\left\lVert u_{1}-u_{2}\right\rVert}{\left\lVert y_{1}-y_{2}\right\rVert}e^{\pm j\angle(u_{1}-u_{2},y_{1}-y_{2})}\mid(y_{1},u_{1}),(y_{2},u_{2})\in R^{-1},u_{1}\neq u_{2},y_{1}\neq y_{2},y_{1},y_{2}\in R(\mathscr{U})\}=\operatorname{SRG}_{R(\mathscr{U})}\setminus\{0,\infty\}$. By [4, p. 588], $0\in\operatorname{SRG}_{\mathscr{U}}(R)\iff\infty\in\operatorname{SRG}_{R(\mathscr{U})}(R^{-1})$ and $\infty\in\operatorname{SRG}_{\mathscr{U}}(R)\iff 0\in\operatorname{SRG}_{R(\mathscr{U})}(R^{-1})$.
  4. d.
    The case $\infty\notin\operatorname{SRG}_{\mathscr{U}}(R)\cup\operatorname{SRG}_{\mathscr{U}}(S)$ is proven entirely analogously to [4, Theorem 6] by taking $\mathrm{dom}(R)=\mathrm{dom}(S)=\mathscr{U}$. If $\operatorname{SRG}_{\mathscr{U}}(R),\operatorname{SRG}_{\mathscr{U}}(S)\neq\emptyset$, the theorem holds for $\infty\in\operatorname{SRG}_{\mathscr{U}}(R)\cup\operatorname{SRG}_{\mathscr{U}}(S)$ by defining +$\infty$+$\infty$=$\infty$ + ∞ = ∞.
  5. e.
[…]
Proof of Theorem 3.
The proof is largely the same as for Theorem 2, where now one input is held fixed. We will only discuss the differences.
  1. a.
    Since $u^{\star}$ is fixed, we cannot use the trick $\frac{\left\lVert R\alpha u-R\alpha u^{\star}\right\rVert}{\left\lVert u-u^{\star}\right\rVert}=\frac{\left\lVert Ru-Ru^{\star}\right\rVert}{\left\lVert(1/\alpha)(u-u^{\star})\right\rVert}$ anymore, since $(1/\alpha)u^{\star}\neq u^{\star}$, unless $u^{\star}=0$ ($\alpha=1$ is trivial). Therefore, the result from Theorem 2 holds without the $R\alpha$ case.
  2. b.
    Since we require that $\left\langle I_{\mathscr{U}}(u-u^{\star}),u-u^{\star}\right\rangle=\left\lVert u-u^{\star}\right\rVert^{2}$, it must hold additionally that $I_{\mathscr{U}}u^{\star}=u^{\star}$.
  3. c.
    Identical to Theorem 2.
  4. d.
    Identical to Theorem 2.
  5. e.
    Identical to Theorem 2, but for the composition of operators one must obey $Ru^{\star}=y^{\star}$.
Proof of Lemma 4.
First we prove the sum rule. By Theorem 2.d., it holds that $\operatorname{SRG}_{\mathscr{U}}(R+S)$ is contained in both sets in the right hand side of Eq. (18), and therefore also in their intersection. Using Theorem 3.d., the same result holds for $\operatorname{SG}_{\mathscr{U},u^{\star}}(R+S)$.
The product rule follows analogously. By Theorem 2.e., it holds that $\operatorname{SRG}_{\mathscr{U}}(TR)$ is contained each of the four sets in the right hand side of Eq. (19), and therefore also in their intersection. Using Theorem 3.e., the same result holds for $\operatorname{SG}_{\mathscr{U},u^{\star}}(TR)$. ∎
Proof of Lemma 5.
Fix $u\in\mathscr{U}$ and define $T_{u}:\mathscr{U}\to\mathscr{U}$ as $T_{u}x:=u-H_{2}H_{1}x$. Note that for all $x,y\in\mathscr{U}$, one has
$\left\lVert T_{u}x-T_{u}y\right\rVert=\left\lVert H_{2}H_{1}x-H_{2}H_{1}y\right\rVert\\ \leq\Gamma(H_{2}H_{1})\left\lVert x-y\right\rVert\leq L\left\lVert x-y\right\rVert,$
where $L=\Gamma(H_{1})\Gamma(H_{2})<1$. By the Banach fixed point Theorem 1, there exists a unique solution $e\in\mathscr{U}$ such that $T_{u}e=e$. Let $y=H_{1}e$, then it is clear that $e=u-H_{2}y=T_{u}e$ holds.
It remains to prove incremental stability. Take $u_{1},u_{2}\in\mathscr{U}$, then by $e_{1}-e_{2}=u_{1}-u_{2}+H_{2}H_{1}e_{2}-H_{2}H_{1}e_{1}$ we obtain $\left\lVert e_{1}-e_{2}\right\rVert=\left\lVert u_{1}-u_{2}\right\rVert+\Gamma(H_{1})\Gamma(H_{2})\left\lVert e_{1}-e_{2}\right\rVert$, hence $\left\lVert e_{1}-e_{2}\right\rVert\leq\frac{1}{1-\Gamma(H_{1})\Gamma(H_{2})}\left\lVert u_{1}-u_{2}\right\rVert$, proving continuity of $e$ in $u$. Continuity of $y$ in $u$ follows from $\left\lVert y_{1}-y_{2}\right\rVert\leq\Gamma(H_{1})\left\lVert e_{1}-e_{2}\right\rVert$, which proves stability. ∎
Proof of Lemma 6.
Almost the same as [19, Theorem 2], but we use the Banach space incremental small gain theorem Lemma 5, instead of the case $\mathscr{U}=\mathscr{Y}=L_{2}$ in [19]. ∎
Proof of Theorem 4.
It is understood that $\operatorname{SRG}(H_{1}):=\operatorname{SRG}_{\mathscr{U}_{p}^{n}}(\tilde{H}_{1})$ and $\operatorname{SRG}(H_{2}):=\operatorname{SRG}_{\mathscr{U}_{q}^{n}}(\tilde{H}_{2})$, where $\tilde{H}_{1}=\iota_{n\leftarrow q}H_{1}\pi_{p\leftarrow n}$ and $\tilde{H}_{2}=\iota_{n\leftarrow p}H_{1}\pi_{q\leftarrow n}$. Therefore, $\tilde{H}_{1}:\mathscr{U}_{p}^{n}\to\mathscr{U}_{q}^{n}$ and $\tilde{H}_{2}:\mathscr{U}_{q}^{n}\to\mathscr{U}_{p}^{n}$ and Eq. (21a) implies $\Gamma(\tilde{H}_{1})<\infty$, $\Gamma(\tilde{H}_{2})<\infty$.
From Eq. (21b) we know that $\mathrm{dist}(\operatorname{SRG}(H_{1})^{-1}+\tau\operatorname{SRG}(H_{2}),0)\geq r$, hence $r_{\mathrm{min}}((\operatorname{SRG}(H_{1})^{-1}+\tau\operatorname{SRG}(H_{2}))^{-1})\leq 1/r$, and so $\Gamma([\tilde{H}_{1},\tau\tilde{H}_{2}])\leq 1/r$.
By Lemma 6, we can conclude that $[\tilde{H}_{1},\tau\tilde{H_{2}}]$ is well-posed for all $\tau\in[0,1]$ and therefore $[\tilde{H}_{1},\tilde{H_{2}}]:\mathscr{U}_{p}^{n}\to\mathscr{U}_{q}^{n}$ with $\Gamma([\tilde{H}_{1},\tilde{H_{2}}])\leq 1/r$.
The final step is to transfer the result to $[H_{1},H_{2}]$. We use the fact that $\pi_{p\leftarrow n}:\mathscr{U}_{p}^{n}\to L_{2}^{p}$ is an isometric isomorphism with inverse $\iota_{n\leftarrow p}$ (by Lemma 1 and 2). For $u,e\in L_{2}^{p},y\in L_{2}^{q}$, define $\tilde{u}=\iota_{n\leftarrow p}u\in\mathscr{U}_{p}^{n},\tilde{e}=\iota_{n\leftarrow p}e\in\mathscr{U}_{p}^{n}$ and $\tilde{y}=\iota_{n\leftarrow q}y\in\mathscr{U}_{q}^{n}$. Since $\iota$ is a bijection, we have
$e=u-H_{2}H_{1}e$ $\iff\tilde{e}=\tilde{u}-\tilde{H}_{2}\tilde{H}_{1}\tilde{e},$
$y=H_{1}e$ $\iff\tilde{y}=\tilde{H}_{1}\tilde{e},$
hence $[H_{1},H_{2}]:L_{2}^{p}\to L_{2}^{q}$ is well-posed if and only if $[\tilde{H}_{1},\tilde{H}_{2}]:\mathscr{U}_{p}^{n}\to\mathscr{U}_{q}^{n}$ is well-posed. Moreover, since $\pi$ is a linear isomorphism, one has $\left\lVert u\right\rVert=\left\lVert\tilde{u}\right\rVert,\left\lVert e\right\rVert=\left\lVert\tilde{e}\right\rVert$ and $\left\lVert y\right\rVert=\left\lVert\tilde{y}\right\rVert$, resulting in
$\Gamma([H_{1},H_{2}])=\Gamma([\tilde{H}_{1},\tilde{H}_{2}]).$
Note that $H_{1}$ and $H_{2}$ are causal if and only if $\tilde{H}_{1}$ and $\tilde{H}_{2}$ are causal. Causality of $[\tilde{H}_{1},\tilde{H}_{2}]$ (and hence $[H_{1},H_{2}]$) follows from applying [29, Theorem 2.11] for the subalgebra of causal operators with finite incremental gain to each application of the Banach fixed point theorem in the proof of 6. ∎
Proof of Lemma 8.
Let $\nu\in[0,1/(\gamma(H_{1})\gamma(H_{2}))$ and write $T_{\nu}=[H_{1},\nu H_{2}]$. By the well-posedness assumption one has $T_{\nu}:L_{2}^{p}\to L_{2\mathrm{e}}^{q}$. By Lemma 7, one has $T_{\nu}:L_{2}^{p}\to L_{2}^{q}$ with $\gamma(T_{\nu})\leq\hat{\gamma}$.
For all $\tau\in[0,1/(\hat{\gamma}\gamma(H_{2}))$, one again applies the small gain theorem to conclude that $T_{\nu+\tau}:L_{2}\to L_{2}$ with $\gamma(T_{\nu+\tau})\leq\hat{\gamma}$. Proceeding inductively $N$ times until $\nu+N\tau=1$, as in the proof of [19, Theorem 2] proves the result. ∎
Proof of Theorem 5.
The proof mimicks the proof of Theorem 4. Note that we assume $H_{1}(0)=0,H_{2}(0)=0$. It is understood that $\operatorname{SG}_{0}(H_{1}):=\operatorname{SG}_{\mathscr{U}_{p}^{n},0}(\tilde{H}_{1})$ and $\operatorname{SG}_{0}(H_{2}):=\operatorname{SG}_{\mathscr{U}_{q}^{n},0}(\tilde{H}_{2})$, where $\tilde{H}_{1}=\iota_{n\leftarrow q}H_{1}\pi_{p\leftarrow n}$ and $\tilde{H}_{2}=\iota_{n\leftarrow p}H_{1}\pi_{q\leftarrow n}$. Therefore, $\tilde{H}_{1}:\mathscr{U}_{p}^{n}\to\mathscr{U}_{q}^{n}$ and $\tilde{H}_{2}:\mathscr{U}_{q}^{n}\to\mathscr{U}_{p}^{n}$ and Eq. (22a) implies $\gamma(\tilde{H}_{1})<\infty$, $\gamma(\tilde{H}_{2})<\infty$.
From Eq. (22b) we know that $\mathrm{dist}(\operatorname{SG}_{0}(H_{1})^{-1}+\tau\operatorname{SG}_{0}(H_{2}),0)\geq r$, hence $r_{\mathrm{min}}((\operatorname{SG}_{0}(H_{1})^{-1}+\tau\operatorname{SG}_{0}(H_{2}))^{-1})\leq 1/r$, and so $\gamma([\tilde{H}_{1},\tau\tilde{H}_{2}])\leq 1/r$.
By Lemma 8 and the well-posedness assumption, we can conclude that $[\tilde{H}_{1},\tilde{H_{2}}]:\mathscr{U}_{p}^{n}\to\mathscr{U}_{q}^{n}$ is well-posed with $\gamma([\tilde{H}_{1},\tilde{H_{2}}])\leq 1/r$.
Since $\mathscr{U}_{p}^{n}$ and $L_{2}^{p}$ ($\mathscr{U}_{q}^{n}$ and $L_{2}^{q}$) are isometrically isomorphic via $\pi_{p\leftarrow n}$ and its inverse $\iota_{n\leftarrow p}$ ($\pi_{q\leftarrow n}$ and $\iota_{n\leftarrow q}$), we can conclude that $[H_{1},H_{2}]:L_{2}^{p}\to L_{2}^{q}$ is well-posed and $\gamma([H_{1},H_{2}])=\gamma([\tilde{H}_{1},\tilde{H}_{2}])$, which proves the claim. ∎
Proof of Theorem 6.
As noted before, the stability of Eq. 11 depends only on the stability of $[\Phi,-G_{\mathrm{zw}}]$, since $G:L_{2}^{p+n_{\mathrm{w}}}\to L_{2}^{q+n_{\mathrm{z}}}$ is assumed to have finite incremental gain. Since $G_{\mathrm{zw}}$ is part of $G$, it follows that $\Gamma(G_{\mathrm{zw}})\leq\Gamma(G)$. If $G_{\mathrm{yw}},G_{\mathrm{zu}}$ are not identically zero, then both $z_{1}\in\operatorname{SRG}(G_{\mathrm{yw}})$ and $z_{2}\in G_{\mathrm{zu}}$ where $z_{1},z_{2}\in\mathbb{C}\setminus\{0\}$. Therefore, the assumption implies that for all $\tau\in[0,1]$
$\mathrm{dist}(\operatorname{SRG}(\Psi)^{-1},\tau\operatorname{SRG}(G_{\mathrm{zw}}))\geq\tilde{r}\geq|z_{1}||z_{2}|/\hat{\Gamma}>0,$ (36)
hence by Theorem 4, $[\Phi,-G_{\mathrm{zw}}]:L_{2}^{n_{\mathrm{z}}}\to L_{2}^{n_{\mathrm{w}}}$ is well-posed with $\Gamma([\Phi,-G_{\mathrm{zw}}])\leq 1/r$, where $r$ is the largest value of $\tilde{r}$ such that Eq. (36) holds.
Now $R=G_{\mathrm{yw}}[\Phi,-G_{\mathrm{zw}}]G_{\mathrm{zu}}+G_{\mathrm{yu}}$ is simply a series and parallel interconnection of operators. Since $[\Phi,-G_{\mathrm{zw}}]:L_{2}^{n_{\mathrm{z}}}\to L_{2}^{n_{\mathrm{w}}}$ is well-posed, we can conclude that $R:L_{2}^{p}\to L_{2}^{q}$ is a well-posed LFR system. From the MIMO SRG definition and Theorem 2 it follows that $\operatorname{SRG}(R)\subseteq\operatorname{SRG}(G_{\mathrm{yu}})+\\ \operatorname{SRG}(G_{\mathrm{yw}})(\operatorname{SRG}(\Phi)^{-1}-\tau\operatorname{SRG}(G_{\mathrm{zw}}))^{-1}\operatorname{SRG}(G_{\mathrm{zu}})$, and therefore $\Gamma(R)\leq\hat{\Gamma}$.
If one of $G_{\mathrm{yw}},G_{\mathrm{zu}}$ is zero, then $R=G_{\mathrm{yu}}$ is well-posed, causal and the gain bound follows from the definition of the SRG.
The causality claim is a result of Theorem 4 and the fact that elements of $RH_{\infty}$ are proper, hence causal. ∎
Proof of Lemma 9.
Let $u_{1},u_{2}\in\mathbb{R}^{d}$ and define $\Delta u=u_{1}-u_{2}$, $\Delta y=y_{1}-y_{2}$ where $y_{1,i}=\phi_{i}(u_{1,i})-\mu u_{1,i}$ and $y_{2,i}=\phi_{i}(u_{2,i})-\mu u_{2,i}$. By [5, Proposition 9], we have
$\Delta u_{i}\Delta y_{i}\geq\frac{1}{\lambda-\mu}\Delta y_{i}^{2},\quad\forall i=1,\dots,d.$ (37)
By summing Eq. (37) and integrating over $\mathbb{R}_{\geq 0}$ one obtains
$\left\langle x-y,\Psi(x)-\Psi(y)\right\rangle\geq\frac{1}{\lambda-\mu}\left\lVert\Psi(x)-\Psi(y)\right\rVert^{2},$ (38)
where $x,y\in L_{2}^{d}$, we substituted $u_{1}=x(t),u_{2}=y(t)$, and $\Psi:=\Phi-\mu I$. By [4, Proposition 2], we conclude $\operatorname{SRG}(\Psi)\subseteq D_{\frac{\lambda-\mu}{2}}(\frac{\lambda-\mu}{2})$, and hence by Theorem 2.a. and 2.b. we conclude $\operatorname{SRG}(\Psi)\subseteq D_{r}(c)$ where $c=\frac{\lambda+\mu}{2}$ and $r=\frac{\lambda-\mu}{2}$. ∎